### v1

# LaT-PFN: A Joint Embedding Predictive Architecture for In-context Time-series Forecasting

Abstract
We introduce LatentTimePFN (LaT-PFN), a foundational Time Series model with a strong embedding space that enables zero-shot forecasting. To achieve this, we perform in-context learning in latent space utilizing a novel integration of the Prior-data Fitted Networks (PFN) and Joint Embedding Predictive Architecture (JEPA) frameworks. We leverage the JEPA framework to create a prediction-optimized latent representation of the underlying stochastic process that generates time series and combines it with contextual learning, using a PFN. Furthermore, we improve on preceding works by utilizing related time series as a context and introducing an abstract time axis. This drastically reduces training time and increases the versatility of the model by allowing any time granularity and forecast horizon. We show that this results in superior zero-shot predictions compared to established baselines. We also demonstrate our latent space produces informative embeddings of both individual time steps and fixed-length summaries of entire series. Finally, we observe the emergence of multi-step patch embeddings without explicit training, suggesting the model actively learns discrete tokens that encode local structures in the data, analogous to vision transformers.
1 Introduction
Time series forecasting is a fundamental task ubiquitous to various domains, ranging from finance [48, 34] to healthcare [32, 28], retail [24, 55], logistics [13] and beyond. Traditional approaches, both statistical [25, 14, 51] and deep-learning-based [47, 49], are frequently incapable of zero-shot forecasting [43] as they require training for each new dataset, limiting generalization capabilities.
In this paper, we introduce LatentTimePFN (LaT-PFN), a foundational Time Series model designed to address zero-shot forecasting by combining the Prior-data Fitted Networks (PFN) and Joint Embedding Predictive Architecture (JEPA) frameworks. LaT-PFN is trained exclusively on a novel synthetic data generation method, allowing us to encode expert knowledge directly in the training data. Furthermore, LaT-PFN learns to perform in-context learning in latent space by approximating the Posterior Predictive Distribution (PPD). We implement this meta-learning approach by enabling the model to learn zero-shot forecasting at test time, with user-provided time series functioning as exemplary context. We train LaT-PFN using an abstract time dimension, which consolidates a wider range of time-frequency patterns, facilitating meta-learning. Additionally, this improves the model’s versatility and accuracy, whilst significantly reducing computational requirements. Following the JEPA methodology, LaT-PFN separates predicting and decoding, which are optimized independently. As such, LaT-PFN is trained to predict the next latent state and exploits a system identification loss as a regularization term, to independently improve the quality of its embedding space.
Our model demonstrates superior zero-shot prediction performance compared to baselines [10, 17, 51, 63], showcasing its effectiveness in handling unseen distributions. Additionally, LaT-PFN produces informative embeddings, demonstrating a comprehensive understanding of time series. Finally, we observe the emergence of multi-step patch embeddings without explicit training, suggesting that LaT-PFN actively learns discrete tokens encoding local structures in the data, reminiscent of vision transformers [18]. Our contributions are as follows:
  • We introduce a novel architecture, LaT-PFN, which introduces a combination of the PFN and JEPA frameworks for zero-shot predictions of univariate time series in latent space.
  • We introduce a novel synthetic prior for simulating context-aware time series data. Specifically, our prior allows for the creation of abstract-time synthetic contexts, defined as collections of time series, similar enough to enable LaT-PFN to approximate the PPD.
  • We demonstrate, through extensive experimentation across a variety of datasets, that LaT-PFN outperforms established baselines with limited training resources, on multiple tasks.
  • Finally, we offer an extensive analysis of the model’s embedding space. This becomes particularly relevant as we notice the emergence of patch-like embeddings, without explicit training. We speculate this may be analogous to Vision Transformers (ViT) [18] and offer evidence it may even amount to a rudimentary time-series-specific corpus.
2 Background
2.1 PFN: Prior-data Fitted Networks
The Prior-data fitted Networks (PFN) methodology [41], is a meta-learning framework that explicitly trains a model for in-context learning. This is achieved through a neural network that approximates the posterior predictive distribution (PPD), thereby approximating Bayesian Inference. This requires a large volume of data, exhibiting a variety of related distributions. To facilitate this, the PFN trains on synthetic data generated by a simulation, whose parameters are sampled from a user-defined prior distribution $Pr(\psi)$. This allows developers to explicitly encode expert knowledge on the family of datasets $\Psi$ that the PFN is trained on, rather than relying on implicitly encoding specialized knowledge in the model’s architecture [41].
The PPD is typically defined as the probability distribution $P(y_{*}|x_{*},\mathcal{D})$, with $\mathcal{D}=\{(x_{i},y_{i})\}^{N}$ being a context dataset. The PFN framework works by approximating this distribution:
$Q_{\theta}(y_{*}|x_{*},\mathcal{D}_{j})\>$ $\approx\>P(y_{*}|x_{*},\mathcal{D})\quad\text{with}\quad\mathcal{D}_{j}\sim S(\mathcal{D}_{j}|\psi_{j}),\quad\psi_{j}\sim Pr(\psi)$
Figure 1: PFN attention [41]
Here $Q$ is a transformer [54] parametrized by $\theta$, $S(.)$ is a simulation engine, and $Pr(.)$ is a prior distribution over instances of the simulation parameters $\psi\in\Psi$. Examples of the relation between features and targets are provided by the tuples $(x_{i},y_{i})$ within $\mathcal{D}$, which are then passed through a transformer encoder [54]. The held-out features $x_{*}$ are passed through the decoder, with a diagonal-only (independent) target mask, allowing only attention over the encoder context (Figure 1). Finally, a head and a cross-entropy loss are applied. The network $Q$ is trained to perform meta-learning on any dataset $\mathcal{D}_{j}$ [41]. The authors prove that this is a valid approximation to the PPD – see Appendix B. Since its introduction, others have applied the PFN framework to various tasks [29, 17, 4, 23]. Notably, TabPFN [29] solves small tabular datasets, zero-shot. Additionally, ForecastPFN [17] introduced this methodology to time-series forecasting thanks to a novel synthetic prior.
2.2 JEPA: Joint Embedding Predictive Architecture
The Joint Embedding Predictive Architecture (JEPA) [35] framework aims to create strong embeddings of any modality, to efficiently handle complex relationships within data. The key concept is to predict state transitions $S_{t}\rightarrow S_{t+1}$ fully in latent space, in order to create embeddings that are not burdened by unnecessary detail. Figure 2 illustrates the core components, which define a sequence of transformations aimed at processing inputs within latent space. Initially, an input embedder $\bar{x}=E_{x}(x)$ converts the input into latent space.
Figure 2: JEPA [35]
This is followed by a predictor $\hat{\bar{y}}=P(\bar{x},z)$ which estimates the next latent state, using the transformed input and a variable $z$, intended to capture relationships not observable in the data space. A target embedder $\bar{y}=E_{y}(x,y)$ then maps the target, taking into account both $x$, $y$ and their the relationship. A criterion, typically a Mean Squared Error (MSE) loss $C(\hat{\bar{y}},\bar{y})$, evaluates the latent prediction. Finally, a generative decoder, $\hat{y}=D(\hat{\bar{y}})$, decodes the latent representations back to the data space.
The JEPA framework has since been applied to image and video data [5, 6]. The target encoder $y_{L}=E_{y}(x,y)$ is updated as an exponential moving average (EMA) of the input encoder $E_{x}$ and followed by a stop gradient, preventing collapse by learning an identity mapping [35, 5]. Other JEPA works include Motion [7], Audio [22], Point-cloud [46] and EEG transfer [27].
3 Methodology
To the best of our knowledge, this work pioneers bringing energy-based JEPA [35] to time series forecasting, as well as combining it with the probabilistic PFN framework [41]. Initially, this integration may appear unconventional; however, we contend that these methodologies are complementary to this modality. JEPA separates concerns of predicting and decoding. This creates an embedding space more suited to the time series modality, which is inherently stochastic. LaT-PFN can thereby explicate latent patterns representing the underlying stochastic process, prioritizing inherent predictability. However, without in-context learning, univariate time series often still lack sufficient predictive power. The context provides samples of the underlying process, enhancing predictability through contextual learning. This is particularly important for time series datasets affected by cold-start and non-stationary issues. For example, one can imagine two series with identical histories but distinct target completions; a context then provides the additional information for extrapolation.
3.1 Problem Statement
Figure 3: Problem statement of a time series forecasting PFN
Starting with a common definition of time series forecasting (TSF), we seek to derive a formulation of TSF for Bayesian Inference. In this work we consider a single supervised TSF problem $P(y|x)$ as predicting future values $y=v_{t+1:H}=[v_{t+1},\dots,v_{H}]$ from timestep $t$ until horizon $H$, conditioned on the history $x=v_{0:t}=[v_{0},\dots,v_{t}]$. It follows that the PPD for this problem is defined as:
$P(y_{*}|x_{*},\mathcal{D})\hskip 1.99997pt=\hskip 1.99997ptP(\>v_{*,t+1:H}\>\>|\>\>v_{*,0:t}\>,\>\>\{(x_{i},y_{i})\}^{N}),\quad(x_{i},y_{i})\hskip 1.99997pt=\hskip 1.99997pt(v_{i,0:t},v_{i,t+1:H})$
Where $\mathcal{D}$ is a context dataset with $N$ examples, for $\mathcal{D}\in\mathbb{R}^{N\times S\times F}$, sequence dimension $S$, example dimension $N$ and features $F$. This differs from the ForecastPFN formulation, which defines the context $\mathcal{D}\in\mathbb{R}^{S\times F}$ as simply the history of the target time series [17]. This interpretation, although elegant, does not allow a context for the TSF task definition and is simply a regular forecasting task with simulated data. See Figure 3 for a visualization.
To complement this notation, let $\bar{x}=\bar{v}_{0:t}$ and latent target $\bar{y}=\bar{v}_{t+1:H}$ be the embedded versions of history $x=v_{0:t}$ and target $y=v_{t+1:H}$ respectively. Similarly, let $\hat{\bar{y}}=\hat{\bar{v}}_{t+1:H}$ be the latent forecast, and $\hat{y}=\hat{v}_{t+1:H}$ be the forecast in data space.
3.2 Abstract Time Dimension
Unlike [17], instead of using absolute time units (year, month, etc.), we map time to a fixed abstract interval $T_{0}:T_{H}$ relative to the forecast start $t$ of each series, with a variable sampling rate – while ensuring the absence of any data leakage. This allows us to map context time series and past segments of the held-out series onto the same interval, independent of their actual place in history. The approach aligns well with the idea of using simulated data to learn Bayesian inference, as it creates a universal and consistent time dimension in which the model can more readily learn general temporal patterns. Furthermore, reducing the number of temporal degrees of freedom that require training, leads to a more efficient convergence and significantly decreases the computational demands for training.
3.3 An Architecture for Latent In-context Forecasting
Figure 4: The LaT-PFN Architecture. The context is embedded in fixed-length series-vectors. These are fed into the PFN Predictor transformer, with the embedded held-out history prompts, using cross-attention. The latent predictions are compared to the latent target, then decoded with a stopgradient, and compared to real targets. Finally, we apply a supervised regularisation on simulation parameters.
Embedder
Figure 5: Production of embeddings and prompts by masking and selection
The embedder $E(.)$ is an eight-layer dilated Mobilenet1D [30]. We empirically found convolutions to be superior to attention over the sequence dimension, in line with [63, 38]. The component is responsible for feature extraction along the sequence dimension. First, we create the context embeddings $\bar{\mathcal{D}}$ and the latent target $\bar{y}$ with time $T$ and value $V$ features. For the latent target specifically, we use an EMA of the embedder parameters, apply a stop gradient, and select only the future timesteps:
$\bar{\mathcal{D}}\hskip 5.0pt$ $=\hskip 5.0ptE_{\theta}(\mathcal{D})$
$\bar{y}\hskip 5.0pt$ $=\hskip 5.0ptSelect_{t+1:H}\big{[}\>StopGrad\big{(}E_{\theta_{\texttt{EMA}}}(x,y)\>\big{)}\>\>\big{]}$
Next, we create the held-out embeddings $\bar{x}$ and the prompt $\bar{z}$:
$\bar{x},\bar{z}\hskip 5.0pt=\hskip 5.0ptE_{\theta}(x)$
To avoid introducing forward-looking values, we apply padding and masking to the input features. The prompt is therefore a function of the historical values and the entire time axis (including the horizon). See Figure 5 for an illustration of the aforementioned process.
Predictor
The predictor $PFN_{\theta}(.)$ forecasts the next latent state, given the prompt and context. We apply learned average pooling $AVG_{\theta}$ over the sequence dimension of the embedded context $\bar{\mathcal{D}}$, by cross-attention with a learned vector $\vec{q}\in\theta$. These are then provided to the encoder of the latent PFN predictor 4, with self-attention over the examples dimension.
$\bar{\eta}\hskip 5.0pt=\hskip 5.0ptAVG_{\theta}(\bar{\mathcal{D}})$
In order to apply 2D attention over the held-out series, we flatten the examples dimension $N$ and sequence dimension $S$ of the prompt $\bar{z}$. Next, we pass it to the decoder, whilst ensuring adjacent sequences from held-out series are not dependent on sharing information.
$M\hskip 3.30002pt=\hskip 3.30002ptdiag(1)^{N\cdot S\times N\cdot S},\quad\hat{\bar{y}}\hskip 3.30002pt=\hskip 3.30002ptPFN_{\theta}(\bar{z},\bar{\eta},M),\quad\mathcal{L}_{\text{latent}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{\bar{y}}-\bar{y})^{2}$
This is achieved by the diagonal-only mask $M$, which we adopt from the original PFN formulation [41]. Consequently, we stay true to the design from Figure 1.
Decoder
The decoder $D_{\theta}(.)$ is a three-layer feedforward network, trained with independent gradients in alignment with the JEPA methodology [35, 5, 6]. We empirically found this to be critical for generating stable embeddings. The decoder utilizes a cross-entropy loss (see Appendix B) to define a categorical distribution over 100 segmented bins of the output space, incorporating label smoothing of 0.01 [50]. This is made feasible by casting the values $V$ to an abstract space through z-normalizing with 2-std. By working in a value-abstract space, the decoder yields an output without making any assumptions on the target distribution – adhering closely to the PFN formulation [41].
$\hat{y}\hskip 3.30002pt=\hskip 3.30002ptD_{\theta}(StopGrad(\hat{\bar{y}})),\quad\mathcal{L}_{\text{decoder}}\hskip 3.30002pt=\hskip 3.30002pt\mathbb{E}_{p(y)}\log p(\hat{y}\>|\>\>\hat{\bar{y}})$
System Identification Head
Inspired by Sim2Real research (e.g. [62]), we add a regularization term, to encourage consistency between history and forecast. This component processes the summarised vectors of both context and predicted held-out series, to identify the simulation parameters that generated the context $\mathcal{D}$. This approach further supervises the latent space, directing it to the underlying characteristics of the simulation, such as trend and seasonality. This, in turn, leads to improved forecasting capabilities and stabilizes training [35]. The latent embeddings and predictions are concatenated and averaged, then passed to a head $H_{\theta}(.)$ alongside the embedded context. $H_{\theta}(.)$ learns a multi-target regression towards unit-normalized simulation parameters.
$\hat{y}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002ptH_{\theta}(AVG_{\theta}(\hat{\bar{y}},\bar{x}),\bar{C}),\quad\mathcal{L}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{y}_{\text{si}}-\psi_{j})^{2}$
Loss
We define our loss as follows:
$\mathcal{L}\hskip 1.99997pt=\hskip 1.99997pt\lambda_{\text{latent}}\mathcal{L}_{\text{latent}}+\lambda_{si}\mathcal{L}_{\text{si}}+\mathcal{L}_{\text{decoder}},\quad\lambda_{\text{latent}}=3.77e^{-3},\quad\lambda_{si}=1e^{-7}$
The decoder is trained following a stop-gradient operation, ensuring one-way independence.
3.4 Context-aware Synthetic Prior with Triple Sampling
(a) ECL
(b) Traffic
(c) Illness
(d) Etth1
(e) Etth2
(f) Synthetic Prior
Figure 6: Decoded fits plotted on the target, with history and target in blue and red scatter respectively, as well as the PPD fore- and backcast mean as blue line with variance, and finally the input example context in green in the background. The fits demonstrate the predictions in latent space have picked up concepts such as seasonality and trend, making it easier to extrapolate after the forecast moment.
Following the PFN example [41, 17, 29], Lat-PFN was trained exclusively on synthetic data. For this work, we adopted and adjusted ForecastPFN’s time series synthetic prior [17]. Like ForecastPFN, we set hyperprior parameters that govern sampling simulation parameters. Unlike ForecastPFN however, our context is defined not by the history of a single time series, but by a collection of example time series. Hence, for most simulation parameters, we introduce a context dimension by employing a three-step sampling strategy:
$\displaystyle\alpha_{c}\hskip 3.00003pt\sim\hskip 3.00003pt$ $\mathcal{U}(\alpha^{*})$
$\mu_{1},\mu_{2}\hskip 3.00003pt\sim\hskip 3.00003pt$ $\mathcal{U}(\alpha_{c}),$
$\psi_{i}\hskip 1.00006pt\sim\hskip 1.99997pt\mathcal{N}(\mu_{1},\sigma^{*}),\quad$ $\psi_{j}\hskip 1.00006pt\sim\hskip 1.99997pt\mathcal{N}(\mu_{2},\sigma^{*}),\quad i,j\in[0,N]$
Specifically, for any simulation parameter in $\psi_{i}$ the procedure is defined as:
  1. 1.
    Hyperprior Parameters $(\alpha^{*})$: Set the global limits for context synthesis.
  2. 2.
    Context Parameters $(\alpha_{c})$: Ranges sampled uniformly from the hyperprior to define the limits within an individual context.
  3. 3.
    Sub-context Parameters $(\mu)$: Cluster centers as $(\mu_{1},\mu_{2})$ sampled uniformly from the context range, are used to introduce contrast within a context and guide the model to recognize when a context example may be non-informative.
  4. 4.
    Series Parameters $(\psi)$: Point estimates normally sampled using a fixed hyperprior variance $\sigma^{*}$, used as simulation parameters in $S(v_{i,0:T}|\psi_{i})$ to generate individual training data series.
This approach ensures balance in inter- and intra-context variance, promoting coherence without trivializing the series and enhancing the model’s ability to generalize in a zero-shot setting. A full description of the synthetic prior and which parameters are used can be found in appendix C.1.
4 Experimental Setup and Training Details
Training
During training, we employ a linear warmup schedule for weight decay, and target encoder ema decay – in line with previous JEPA work [5, 6]. We trained the final model for 24 hours on a single NVIDIA A10G Tensor Core GPU – per seed. Additionally, we employ tuning on schedule parameters and $\mu$-parametrization, which was essential for training stability [59]. For more training and tuning details, see Appendix C.
Metrics
We report the mean and error statistics of the zero-shot performance of multiple seeds. These are accuracy, MSE per time-step, and the cumulative relative root mean squared error (CRRMSE). The latter is normalized by the cumulative target. We motivate choosing CRRMSE for being scale-invariant, and practical applicability in tasks such as sales forecasting.
Benchmarks & Context Curation
For real datasets, fixed sequence length contexts were provided to the model. These were mapped to an abstract interval and Z-normalized with 2-std using the history. Which samples to include as contexts is coined Context Curation, and is a process analogous to Prompt Engineering in Large Language Models (LLMs) [40, 56]. For details, see Appendix C.
Baselines
We selected FBProphet [51], ARIMA [10], and ForecastPFN [17] as forecasting baselines, all of which have open-source codebases. We compared against TS2Vec, in a downstream classification task, [63] to evaluate the quality of the embedding space. We use the original authors’ codebase and either download the trained model weights or re-train, wherever applicable. Finally, all baselines are tuned, where applicable, and then evaluated on identical datasets. For more details on baselines, see Appendix C.
Table 1: Forecasting scores (with std) for different time horizons. Results demonstrate a strong performance for all of the benchmarks from decoded latent predictions. For details see Appendix C
Mean Squared Error (MSE)
ECL ETTh1 ETTh2 Illness Traffic
Arima 0.44 0.44 0.45 0.50 0.25
ForecastPFN 1.16 0.84 0.74 3.05 0.24
FBProphet 0.61 0.98 0.67 0.96 0.29
LaT-PFN 0.32 $\pm$ 4.0e-2 0.59 $\pm$ 3.0e-2 0.45 $\pm$ 1.8e-2 0.23 $\pm$ 2.3e-2 0.23 $\pm$ 2.8e-2
Cumulative Relative Root Mean Squared Error (CRRMSE)
Arima 5.83 4.93 4.74 5.02 2.68
ForecastPFN 11.39 7.07 6.66 15.6 4
FBProphet 5.75 8.33 6.42 7.79 3.34
LaT-PFN 4.03 $\pm$ 2.4e-1 4.75 $\pm$ 2.1e-1 3.95 $\pm$ 1.5e-1 3.11 $\pm$ 1.8e-1 2.01 $\pm$ 2.0e-1
5 Results & Discussion
Forecasting
Table 2: Overall accuracy of UCR datasets, by fitting SVM on top of fixed-length summary embeddings [20]. Per-dataset scores are in Appendix E.
training budget accuracy
TS2Vec finetuned 45.7%
LaT-PFN zero-shot 51.1%
We find that LaT-PFN exhibits strong, zero-shot forecasting performance across a wide variety of datasets, containing very different temporal patterns and levels of noise. We outline the results of the zero-shot forecasting experiments in Table 1. As can be observed, in nearly all cases LaT-PFN outperforms the baselines. Furthermore, we demonstrate visual curve-fits in Figure 6. Here we can see the model understands useful patterns from the context, such as trend and seasonality, and uses them to reach a good extrapolation for the held-out targets. This suggests that LaT-PFN is inherently strong at generalizing across distributions for zero-shot forecasting, which makes it widely applicable for downstream tasks.
Figure 7: T-SNE [53] of fixed-length embeddings by dataset
Next, we investigate the importance of the context by evaluating the uplift in performance on a synthetic validation set, for different context sizes. We chose to carry out this experiment on simulated data to guarantee a control environment. The result is visualized in Figure 8, which shows a clear correlation between performance and context size. More interestingly, this pattern persists even beyond the original context size of the model during training (represented by the vertical dotted line). This validates our assumption that the key to carrying out in-context learning for zero-shot forecasting is to provide a rich and substantial set of examples. This can be especially impactful when applied in a production setting with an abundance of example series, such as sales forecasting.
Figure 8: Extrapolating evaluation context generalizes beyond training context size
Classification
We evaluate the quality of LaT-PFN’s embedding space by introducing a second downstream task, namely time series classification on the 129 multi-domain UCR datasets. We compare our model to TS2Vec, a state-of-the-art universal representation learning framework for time series [63]. We train TS2Vec on the UCR datasets, then fit Support Vector Machines (SVM) [8] on their embedding space. We compare this to an SVM fitted on the fixed-length summary embeddings of a frozen LaT-PFN, trained on synthetic data only. As is evident in the results reported in Table 2, LaT-PFN outperforms TS2Vec across the entire benchmark of datasets. See Appendix E for individual dataset scores.
We argue these results may be due to bias introduced by time series-specific heuristics and contrastive positive pair selection, both of which TS2Vec and similar methods rely heavily upon [63, 60]. By contrast, LaT-PFN independently discovers the most representative embedding space in which self-supervised learning is most optimal. Furthermore, TS2Vec is trained on the datasets it is evaluated on. This may traditionally be thought of as an advantage. Yet, often these datasets are small and lack variety, increasing the risk of out-of-distribution data in the test set. On the other hand, LatPFN’s prior ensures a wider variety of contexts, whilst adapting to the specific context during test time.
Exploring the Embedding Space
Figure 10: Found patches (red) and top-3 closest matches (blue)
We carry out a qualitative analysis of the latent space generated by LaT-PFN. When analyzing the fixed-length summary embeddings after dimensionality reduction, we observe a clear separation in clusters by dataset – see Figure 7. This suggests that LaT-PFN picks up distributional differences between these domains, making these fixed-length summary embeddings useful for search, classification and other downstream applications.
Figure 9: Embeddings per-timestep show patch-like behavior. More in Appendix D
Interestingly we can observe the emergence of distinct, regular, and discrete patch-like tokens in the heatmaps of the embedding space. An example of this is visualized in Figure 9. Whilst not always humanly interpretable, discernible patterns often emerge from a closer analysis of these patches. For instance, Figure 10 plots some examples in the data space, alongside their closest matches when taking the $L_{2}$-distance between the average patch-embedding. We can observe that these are visually related and seem to describe specific local features. This presents similarities to the patch-like processing of visual "tokens" in Vision Transformers (ViT) and their approach to high-level feature representation [18]. LaT-PFN, however, appears to have learned these tokens independently, without the need to outright encode this pattern in the model’s architecture.
Finally, after applying dimensionality reduction to the set of patch embeddings (Figure 11), we notice the clear emergence of a few clusters. Given the aforementioned similarity between shapes of patches in data space with similar token embeddings, we speculate these could be a latent vocabulary that has been independently learned by the model, consisting of common "words" amongst multi-domain series. This would once more validate our core assumption that, despite time series data often being perceived as highly domain-dependant, there is in fact enough commonality to warrant the use of meta-learning approaches to create foundational models for this data modality.
6 Related Work
Figure 11: DBScan [19] clustering of patches.
For time series forecasting, focus has recently shifted from statistical models such as ARIMA [10], ES [25] and FBProphet [51] towards deep learning. Early pioneers are DeepAR [47] and hybrid methods [49]. Subsequently, N-Beats [42], N-hits [12], Informer [65] and FEDformer [66], combine classical with deep learning paradigms.
More recently, following the success of the NLP domain, many attempts have been made to leverage the transformer architecture [54] to create foundation models [9] for time series forecasting. The two main approaches for doing so can be categorized as; (a) re-purposing LLMs [26, 67, 31] and (b) training transformer models from scratch [17, 4, 57, 15, 61, 45]. By contrast, our research seeks to integrate representation- and in-context learning in time series forecasting, combining strong performance with a low computational budget.
Another approach altogether is to create embeddings and leave the downstream application to the discretion of end users. Notable are TS2Vec [63], TimeCLR [60], SimMTM [16] and TF-C [64]. These methods often rely heavily on expert heuristics, whereas our approach serves both as an embedder and a forecaster, instead leveraging a self-supervised latent space to drive predictions.
7 Limitations
The research presented has a few limitations. Firstly, the current framework is constrained to univariate series, limiting its applicability to multivariate scenarios. Similarly, this model has yet to prove effective in handling hierarchical, discrete, or zero-inflated series, which are often underrepresented in this research field. Defining appropriate synthetic priors could address this, which we leave to future work. Secondly, we observe a low level of standardization for time series deep learning research. Whilst we found an (over) abundance of standard datasets, we observed that data processing, and target and horizon selection are still arbitrary, reducing the comparability of results across studies. This is beyond the scope of this paper but may be addressed in future work. Finally, we designed our model to be highly adaptive to a given context. As a result, sub-optimal contexts can lead to variability in performance. We argue this is analogous to prompt engineering, which assigns this responsibility to the expert user. In future work, we hope to address this by exploring automated methods – via retrieval augmented generation (RAG) [37] or prompt tuning [36]. Addressing these points will be valuable for enhancing the model’s versatility.
8 Conclusion
In this paper, we proposed LaT-PFN, a novel architecture for zero-shot univariate time series forecasting. The model integrates both the PFN and JEPA architecture to allow for in-context learning in latent space. LaT-PFN is trained exclusively on synthetic data, for which we define a novel prior. Combined with an abstract time dimension, this makes our model extremely versatile and able to achieve strong forecasting results across different data distributions, time granularities, and forecast horizons. Finally, its strong embedding space exhibits interesting emergent qualities, such as the development of patch-like tokens that resemble a corpus of time series features. This makes LaT-PFN ideal for supporting transfer learning on downstream models.
Acknowledgments and Disclosure of Funding
This research was fully funded by WAIR (AI4R B.V.). We would like to extend our sincere gratitude to our team and all contributors whose invaluable assistance and insightful suggestions made this work possible.
Appendix / supplemental material
Appendix A Broader Societal Impacts
It is our belief that the work presented in this paper does not pose major negative societal risks. The model is trained exclusively on synthetic time series data, thus eliminating any risk related to data privacy, disinformation or consent. On the contrary, given how ubiquitous time series are in the real world, we hope that our model for zero-shot forecasting will provide a net benefit for society, for a few different reasons. Firstly, it is significantly more efficient to train than many of the baselines we evaluate against, as pointed out in Section 4. Secondly, as a zero-shot foundation model, it only requires training once, before being used out-of-the-box. Thirdly, many sectors in society could benefit from having improved forecasting abilities for assisting day-to-day decision-making. We hope this will result in widespread adoption of pre-trained models for forecasting tasks, thus improving industry effectiveness and reducing energy consumption and CO2 emissions.
Appendix B Derivation PFN NLLLoss
This derivation shows how to approximate Bayesian Inference via a PFN. It is adapted from [41] and forms the justification of why using a simulation with a cross-entropy loss is an approximation of the PPD. Assume we want to approximate PPD $P(y|x,\mathcal{D})$ with network $Q(y|x,\mathcal{D})$. We start with the KL-divergence between the two and end up at entropy function $\bm{H}()$ and constant $C$. The latter does not depend on parameters $\theta$ and is dropped in optimisation.
[…]
We can then integrate the PFN simulation by sampling the data from our prior:
[…]
Appendix C Reproducibility Details
This section is dedicated to providing comprehensive reproducibility details for the research presented in this document. By including this appendix, we aim to uphold transparency standards in research.
C.1 Exhaustive Description Synthesis Training Data
Following the example set by [17] we define a synthetic prior that leverages the underlying components of time series data, namely trend, seasonality, and noise. The trend parameter is made up of linear and exponential components; the seasonality parameter is made up of annual, monthly, and weekly components and the noise is derived from a Weibull distribution:
$y_{t}=\psi(t)\cdot z_{t}=trend(t)\cdot seasonality(t)\cdot z_{t}$
The linear and exponential trend components are defined by two constituent parts each: scaler and offset. The scaler $m$ is a context-wide parameter, therefore we use triple sampling to generate them. The offset $c$ is domain-specific, so we directly sample them, uniformly, from a range of hyperpriors. The formula for the trend component is the same as defined by [17]:
$trend(t)=1+(m_{lin}\cdot t+c_{lin})\cdot(m_{exp}\cdot c_{t}^{exp})$
The seasonality parameters are all context-dependant, therefore we apply triple sampling to derive them. Additionally, we introduce a new parameter, $\delta$ to define the number of frequency features for each seasonality parameter. We use triple sampling for the seasonality parameters and direct univariate sampling for the $\delta$ parameter. The formulas have once again been adapted from [17]:
$seasonal(t)=seasonal_{week}(t)\cdot seasonal_{month}(t)\cdot seasonal_{year}(t)$
$\delta_{v}\sim\mathcal{U}(\alpha^{*}_{\delta})$
$seasonal_{v}(t)=1+m_{v}\sum_{f=1}^{\delta_{v}}\left[c_{f,v}\sin\left({2f\pi\frac{t}{p_{v}}}\right)+d_{f,v}\cos\left({2f\pi\frac{t}{p_{v}}}\right)\right]$
$v\in{\{week,month,year\}}$
$p_{week}=7,~{}p_{month}=30.417,~{}p_{year}=30.417$
$c_{v}\sim\mathcal{N}\left(0,\frac{1}{\lvert\delta_{v}\rvert}\right),~{}d_{v}\sim\mathcal{N}\left(0,\frac{1}{\lvert\delta_{v}\rvert}\right)$
The noise component follows the definition laid out by [17], the only difference being that we sample a single $k$ parameter per context, and use it in the Weibull distribution:
$z=1+m_{noise}(z-\overline{z})$
$z\sim Weibull(1,k),~{}\overline{z}=\left(\ln{2}\right)^{1/k},~{}m_{noise}\in{\{\mathcal{U}(0,0.1),\mathcal{U}(0.2,0.4),\mathcal{U}(0.6,0.8)\}}$
Unlike [17], we train using an abstract time dimension. In order to include many different types of temporal patterns, we define a resolution parameter $\rho$, which determines how many points are represented per interval, for each individual time series. We define this as the relation between the sequence dimension and resolution parameter:
$t=\frac{S}{\rho}\$
Finally, for the resolution and exponential-trend parameters, we obtain the context parameters $\alpha_{c}$, described in Section 3.4, by sampling log-uniformly, due to the logarithmic nature of these components. Specifically, we guarantee a median sample value without having a forced symmetric range around that median value. To do so, we (a) define a mapping to log space, (b) sample uniformly, and (c) inverse the mapping after. The formula is as follows:
$\displaystyle map(x,\kappa)\quad$ $=\quad log_{2}(x\cdot\kappa+1)$
$map^{-1}(x,\kappa)\quad$ $=\quad\frac{2^{x}-1}{\kappa}$
$\alpha_{c,mapped}\quad$ $\sim\quad\mathcal{U}\Big{(}map(\alpha^{*},\kappa)\Big{)}$
$\alpha_{c}\quad$ $=\quad map^{-1}(\alpha_{c,mapped},\kappa)$
For multiplier $\kappa$ and function $map(.)$. Please see Table 3 for details on the multipliers and hyperpriors.
Table 3: Hyperpriors for Simulation Engine
Hyperparameter Min Value Max Value
Seasonality
Annual Frequency Scale $-8.0$ $8.0$
Monthly Frequency Scale $-4.0$ $4.0$
Weekly Frequency Scale $-2.0$ $2.0$
Variance 0.15
Trend
Linear $-0.015$ $0.015$
Linear Variance 0.005
Exponential $0.996$ $1.0016$
Exponential Variance 0.001
Exponential Multiplier $\kappa_{m_{exp}}$ 507
Noise
Noise ($k$) $0.8$ $5$
Resolution
Resolution Min $0.1$ $1.0$
Resolution Multiplier $\kappa_{\rho}$ 53.6
Offset
Linear Offset $-1$ $2$
Exponential Offset $-1$ $2$
Harmonics
$\delta_{\text{harmonics}}$ $4$ $12$
C.2 Baselines
Forecasting
For forecasting we use the following baselines:
  • FBProphet [51], a common industry standard with respect to the type of univariate series we model. We use the publicly available repository [21] and manually tune the parameter changepoint_prior_scale=1.0.
  • ARIMA [14], another classic baseline from the industry. We use the Statsmodel implementation [44] publicly available and pass the argument order=(5, 1, 0).
  • ForecastPFN [17] is a flexible in-context forcaster model and one of the inspirations for this work. We download their open-source repository available under Apache Licence 2.0 and trained model weights. We adopt their inference script to work in our repository. We have tried several data normalizations by hand to get the most out their model and picked the best one, which turned out to be Z-normalization with 2std – on top of the normalizations present in the repository already.
Classification
For classification we use the TS2Vec baseline [63], as it is a strong multi-purpose representation learner and also an inspiration for this work, although TS2Vec does not claim to implement zero-shot forecasting. They open-source their training code under the MIT Licence, so we train TS2Vec from scratch on the train splits of the UCR [20] archive, after which we evaluate on the corresponding test splits with the resulting frozen TS2Vec architecture
C.3 Benchmark Preprocessing
We use five datasets for our experiments, detailed in section 4. These are considered standard datasets for benchmarking the performance of time series forecasting models [43, 58, 65].
For each dataset, we create univariate time series out of each of the columns. Additionally, we further split up the univariate series along the time dimension to create single entities of different granularity, which we then use for context creation. We normalize the time dimension of each dataset so that it fits in the abstract $[-3,1]$ interval. Furthermore, we apply Z-score normalization, with 2 standard deviations, to the values of each univariate time series.
The specifics of each dataset are presented in the following sections. We reference the original sources of the datasets. However, practically we obtained the actual files from the ForecastPFN repository [17, 2]
C.3.1 Illness
Illness is a dataset of influenza-like illness patients in the United States. It reports patients data in a weekly time granularity, from 1997 to 2024. [1]
During pre-processing, we create files for each column of the dataset and split them along the time dimension, with a monthly periodicity.
We define a sequence length of 160 equidistant intervals, of which the last 25% are used as the prediction target. We use a rolling window over the dataset, with a stride of 1.
We hand-picked 4 windows of the historic series as context examples and kept 1 as held-out starting from forecast date 01-01-2018, which we evaluate the zero-shot performance on.
Context Held-out
ILITOTAL 2008-2012 ILITOTAL 2018-2022
ILITOTAL 2013-2017
TOTAL_PATIENTS 2008-2012
TOTAL_PATIENTS 2013-2017
C.3.2 EttH1 & EttH2
The ETT (Electricity Transformer Temperature) dataset is an electricity power deployment dataset. The data is reported both hourly and minute-wise, from 2016-07-01 to 2018-06-26. The dataset reports two different series, ETT1, and ETT2, corresponding to two different transformer stations, where the data was recorded [65].
We experiment on the hourly series but aggregate the data monthly and create univariate time series out of each column. Following the dataset’s original directions, we use the OT column as held-out, on which we evaluate the zero-shot performance of the model.
We define a sequence length of 240 equidistant intervals, of which the last 25% are used as targets. We use a rolling window over each time series with a stride of 10.
We run 2 separate experiments for this dataset, one for each transformer station series. We change both the held-out series and context for each experiment.
  • Experiment 1: ETTh1
    We provide a context dimension of 14 and leave one held-out time series.
    Context Held-out
    ETTh1 HUFL 2017/01 - 2017/05 ETTh1 OT 2018/01 - 2018/05
    ETTh1 HULL 2017/01 - 2017/05
    ETTh1 LUFL 2017/01 - 2017/05
    ETTh1 LULL 2017/01 - 2017/05
    ETTh1 MUFL 2017/01 - 2017/05
    ETTh1 MULL 2017/01 - 2017/05
    ETTh1 OT 2017/01 - 2017/05
    ETTh2 HULL 2017/01 - 2017/05
    ETTh2 LUFL 2017/01 - 2017/05
    ETTh2 LULL 2017/01 - 2017/05
    ETTh2 MUFL 2017/01 - 2017/05
    ETTh2 MULL 2017/01 - 2017/05
    ETTh2 OT 2017/01 - 2017/05
  • Experiment 2 - ETTh2:
    We provide a context dimension of 4 and leave one held-out time series.
    Context Held-out
    ETTh1 HUFL 2017/01 - 2017/05 ETTh2 OT 2018/01 - 2018/05
    ETTh1 OT 2017/01 - 2017/05
    ETTh2 HUFL 2017/01 - 2017/05
    ETTh2 OT 2017/01 - 2017/05
C.3.3 Traffic
Traffic is a dataset reporting freeway congestion in California, USA. We used the data from [17], who originally sourced it from [11]. The data is aggregated daily and consists of 860 numerical columns (named 1-860), plus a target column (OT).
We aggregate the data monthly and create univariate time series for each of the columns.
We define a sequence length of 240 equidistant intervals, of which the last 25% are used as targets. We use a rolling window over each time series with a stride of 1.
We provide a context dimension of 8 and leave two held-out time series.
Context Held-out
360 2016/11 - 2017/05 360 2017/11 - 2018/05
360 2016/10 - 2017/04 360 2017/08 - 2018/02
293 2016/11 - 2017/05
293 2016/10 - 2017/04
710 2016/11 - 2017/05
710 2016/10 - 2017/04
405 2016/11 - 2017/05
405 2016/10 - 2017/04
C.3.4 ECL
ECL (Electricity Consuming Load) [52] reports electricity consumption in Kwh with an hourly frequency from 2016-07-01 to 2019-07-02. The data consists of 320 numerical columns (named 0-319), plus a target column (OT)
We aggregate the data daily and create univariate time series for each of the columns.
We define a sequence length of 240 equidistant intervals, of which the last 25% are used as targets. We use a rolling window over each time series with a stride of 10.
We provide a context dimension of 15 and leave two held-out time series.
Context Held-out
OT 2019/01/01 - 2018/01/14 OT 2019/05/01 - 2018/05/14
OT 2019/01/07 - 2018/01/20 OT 2019/04/01 - 2018/04/14
OT 2019/01/17 - 2018/01/30
OT 2019/02/01 - 2019/02/14
OT 2019/02/07 - 2019/02/20
OT 2019/02/17 - 2019/03/02
OT 2019/03/01 - 2019/03/14
OT 2019/03/07 - 2019/03/20
OT 2019/03/17 - 2019/03/30
OT 2017/04/01 - 2017/04/14
OT 2017/04/07 - 2017/04/20
OT 2017/04/17 - 2017/04/30
C.4 Training & Tuning
Please consider the following training and tuning details:
  • We use a Cosine Annealing With Decay and Warm restarts Learning rate schedule. Base learning rate is 9e-4 and the scheduler parameters consist of a decay of 0.96 and a T0 of 9.
  • We train with a linear warmup of 95 epochs for JEPA-EMA decay and weight decay. Increasing the former from 0.9952 to 1.0 and the latter from 1.77e-4 to 4.9e-2.
  • We tune all of the above hyperparameters with Optuna [3] in the 3 runs, spanning a total of 216 hours.
  • We apply MUP [59], which allows us to tune a smaller model and zero-shot transfer the hyperparameters we found. This also proved essential for training stability. MUP also provides us with the MUP-AdamW optimizer [59, 39, 33], which was the optimizer in all experiments.
  • All the variables in the bullets below were not tuned but hand-picked by either trial and error or expert knowledge.
  • We apply TF32 quantization but no mixed-precision training since this gave stability issues
  • For the final model we train with a batch size of 32, a context size of 14, a history of 180, horizon of 60 and held-out size of 2.
  • For system identification we multi-target regress to 10 simulation parameters, unit-normalized. Specifically, these variables and normalization scales are:
    • Annual frequency scale: unit-normalized
    • Monthly frequency scale unit-normalized
    • Weekly frequency scale: unit-normalized
    • Trend linear: unit-normalized
    • Trend exponential: log-unit-normalized
    • Offset linear: unit-normalized
    • Offset exponential: unit-normalized
    • Noise scale: log-unit-normalized
    • Resolution: log-unit-normalized
  • Since we have a synthetic dataset, we have no natural end of an epoch. We elect to choose a 250-batch epoch before updating schedules.
  • We map our training data to an abstract time domain of $[-3,1]$, with the forecast moment at the origin.
  • We define 100 bins on the abstract $[-3.5,3.5]$ range, as the output space for the decoder.
  • Model-width is set at 512 neurons for the final model and 128 neurons for tuning. Corresponding depths are:
    • Embedder: 8 layers
    • Predictor: 3 layers
    • SI-head: 2 layers
    • Decoder: 3 layers
  • On top of the compute documented in Section 4, which indicates the compute required to train one seed of the final model, we additionally used more compute for experiments in the discovery phase. How much is unfortunately not something we have kept track of.
Figure 12: PCA plot of fixed-length summary embeddings grouped by domain
Appendix D Latent Space Visualised
(a) ETTh1
(b) ETTh2
(c) Traffic
Figure 13: Additional per-step time series embeddings
Appendix E Additional results
E.1 UCR [20] Accuracy by Dataset
Table 4: UCR scores per dataset: A-F
UCR Datasets Our work TS2Vec
ACSF1_TEST 37.00% 41.00%
Adiac_TEST 2.05% 2.05%
AllGestureWiimoteX_TEST 29.14% 25.14%
AllGestureWiimoteY_TEST 25.00% 17.86%
AllGestureWiimoteZ_TEST 22.86% 16.43%
ArrowHead_TEST 46.29% 41.71%
BME_TEST 54.00% 46.67%
Beef_TEST 50.00% 40.00%
BeetleFly_TEST 60.00% 85.00%
BirdChicken_TEST 60.00% 75.00%
CBF_TEST 53.33% 47.56%
Car_TEST 21.67% 21.67%
Chinatown_TEST 60.06% 76.38%
ChlorineConcentration_TEST 53.26% 53.26%
CinCECGTorso_TEST 29.06% 25.58%
Coffee_TEST 75.00% 82.14%
Computers_TEST 56.40% 66.40%
CricketX_TEST 29.23% 13.85%
CricketY_TEST 24.87% 17.44%
CricketZ_TEST 29.49% 14.36%
Crop_TEST 38.66% 32.28%
DiatomSizeReduction_TEST 30.07% 30.07%
DistalPhalanxOutlineAgeGroup_TEST 67.63% 46.76%
DistalPhalanxOutlineCorrect_TEST 58.33% 58.33%
DistalPhalanxTW_TEST 58.27% 30.22%
DodgerLoopDay_TEST 23.75% 37.50%
DodgerLoopGame_TEST 56.52% 50.00%
DodgerLoopWeekend_TEST 87.68% 89.13%
ECG200_TEST 64.00% 64.00%
ECG5000_TEST 82.69% 87.51%
ECGFiveDays_TEST 49.71% 49.71%
EOGHorizontalSignal_TEST 17.13% 29.83%
EOGVerticalSignal_TEST 21.82% 20.99%
Earthquakes_TEST 74.82% 74.82%
ElectricDevices_TEST 51.87% 54.30%
EthanolLevel_TEST 25.40% 26.40%
FaceAll_TEST 26.63% 42.43%
FaceFour_TEST 44.32% 18.18%
FacesUCR_TEST 27.46% 18.29%
FiftyWords_TEST 22.64% 12.53%
Fish_TEST 14.29% 22.29%
FordA_TEST 76.52% 61.52%
FordB_TEST 65.43% 53.95%
FreezerRegularTrain_TEST 76.70% 75.75%
FreezerSmallTrain_TEST 76.00% 75.86%
Fungi_TEST 60.22% 46.24%
Table 5: UCR scores per dataset: G-R
UCR Datasets Our work TS2Vec
GestureMidAirD1_TEST 20.77% 20.00%
GestureMidAirD2_TEST 26.92% 20.77%
GestureMidAirD3_TEST 16.15% 10.00%
GesturePebbleZ1_TEST 40.12% 50.58%
GesturePebbleZ2_TEST 42.41% 31.01%
GunPointAgeSpan_TEST 58.23% 57.91%
GunPointMaleVersusFemale_TEST 62.66% 84.81%
GunPointOldVersusYoung_TEST 52.38% 100.00%
GunPoint_TEST 49.33% 64.67%
Ham_TEST 59.05% 51.43%
HandOutlines_TEST 64.05% 64.05%
Haptics_TEST 22.40% 20.78%
Herring_TEST 59.38% 59.38%
HouseTwenty_TEST 69.75% 71.43%
InlineSkate_TEST 18.18% 18.73%
InsectEPGRegularTrain_TEST 100.00% 100.00%
InsectEPGSmallTrain_TEST 100.00% 100.00%
InsectWingbeatSound_TEST 30.15% 21.26%
ItalyPowerDemand_TEST 61.22% 49.85%
LargeKitchenAppliances_TEST 56.27% 57.60%
Lightning2_TEST 54.10% 54.10%
Lightning7_TEST 30.14% 41.10%
Mallat_TEST 25.54% 12.32%
Meat_TEST 68.33% 53.33%
MedicalImages_TEST 51.45% 51.45%
MelbournePedestrian_TEST 24.52% 62.03%
MiddlePhalanxOutlineAgeGroup_TEST 45.45% 18.83%
MiddlePhalanxOutlineCorrect_TEST 57.04% 57.04%
MiddlePhalanxTW_TEST 46.10% 27.27%
MixedShapesRegularTrain_TEST 59.13% 52.21%
MixedShapesSmallTrain_TEST 48.78% 48.49%
MoteStrain_TEST 87.22% 77.32%
NonInvasiveFetalECGThorax1_TEST 21.63% 1.83%
NonInvasiveFetalECGThorax2_TEST 26.97% 1.83%
OSULeaf_TEST 35.95% 18.18%
OliveOil_TEST 40.00% 40.00%
Overall 51.20% 45.74%
PLAID_TEST 31.10% 24.39%
PhalangesOutlinesCorrect_TEST 61.31% 61.31%
Phoneme_TEST 15.98% 11.29%
PickupGestureWiimoteZ_TEST 38.00% 52.00%
PigAirwayPressure_TEST 5.77% 19.71%
PigArtPressure_TEST 4.81% 25.96%
PigCVP_TEST 5.29% 46.63%
Plane_TEST 22.86% 9.52%
PowerCons_TEST 87.22% 85.56%
ProximalPhalanxOutlineAgeGroup_TEST 82.44% 48.78%
ProximalPhalanxOutlineCorrect_TEST 68.38% 68.38%
ProximalPhalanxTW_TEST 63.90% 35.12%
RefrigerationDevices_TEST 39.20% 54.40%
Rock_TEST 46.00% 34.00%
Table 6: UCR scores per dataset: S-Z
UCR Datasets Our work TS2Vec
ScreenType_TEST 45.33% 40.80%
SemgHandGenderCh2_TEST 74.33% 58.83%
SemgHandMovementCh2_TEST 30.89% 35.11%
SemgHandSubjectCh2_TEST 37.11% 35.56%
ShakeGestureWiimoteZ_TEST 54.00% 64.00%
ShapeletSim_TEST 50.56% 87.78%
ShapesAll_TEST 28.00% 42.67%
SmallKitchenAppliances_TEST 40.53% 55.47%
SmoothSubspace_TEST 57.33% 78.00%
SonyAIBORobotSurface1_TEST 59.73% 42.93%
SonyAIBORobotSurface2_TEST 78.07% 61.70%
StarLightCurves_TEST 82.93% 84.74%
Strawberry_TEST 64.32% 64.32%
SwedishLeaf_TEST 31.68% 8.48%
Symbols_TEST 55.78% 17.39%
SyntheticControl_TEST 50.33% 77.33%
ToeSegmentation1_TEST 70.61% 65.79%
ToeSegmentation2_TEST 80.77% 80.77%
Trace_TEST 65.00% 19.00%
TwoLeadECG_TEST 49.96% 49.96%
TwoPatterns_TEST 31.62% 25.87%
UMD_TEST 49.31% 50.00%
UWaveGestureLibraryAll_TEST 51.84% 12.37%
UWaveGestureLibraryX_TEST 57.31% 22.28%
UWaveGestureLibraryY_TEST 44.89% 20.38%
UWaveGestureLibraryZ_TEST 52.07% 31.38%
Wafer_TEST 89.21% 89.21%
Wine_TEST 50.00% 50.00%
WordSynonyms_TEST 29.31% 21.94%
WormsTwoClass_TEST 62.34% 57.14%
Worms_TEST 46.75% 42.86%
Yoga_TEST 52.97% 53.57%
End of the Appendix.

### v2

@@
 We leverage the JEPA framework to create a prediction-optimized latent representation of the underlying stochastic process that generates time series and combines it with contextual learning, using a PFN.
-Furthermore, we improve on preceding works by utilizing related time series as a context and introducing an abstract time axis.
-This drastically reduces training time and increases the versatility of the model by allowing any time granularity and forecast horizon.
+Furthermore, we improve on preceding works by utilizing related time series as a context and introducing a normalized abstract time axis.
+This reduces training time and increases the versatility of the model by allowing any time granularity and forecast horizon.
 We show that this results in superior zero-shot predictions compared to established baselines.
@@
 We implement this meta-learning approach by enabling the model to learn zero-shot forecasting at test time, with user-provided time series functioning as exemplary context.
-We train LaT-PFN using an abstract time dimension, which consolidates a wider range of time-frequency patterns, facilitating meta-learning.
-Additionally, this improves the model’s versatility and accuracy, whilst significantly reducing computational requirements.
+We train LaT-PFN using a normalized time-axis, which consolidates a wider range of time-frequency patterns, facilitating meta-learning.
+Additionally, this improves the model’s versatility and accuracy, whilst reducing computational requirements.
 Following the JEPA methodology, LaT-PFN separates predicting and decoding, which are optimized independently.
@@
 Our contributions are as follows:
-  • We introduce a novel architecture, LaT-PFN, which introduces a combination of the PFN and JEPA frameworks for zero-shot predictions of univariate time series in latent space.
+  • We introduce a novel architecture, LaT-PFN, which implements a combination of the PFN and JEPA frameworks for zero-shot predictions of univariate time series in latent space.
   • We introduce a novel synthetic prior for simulating context-aware time series data.
-Specifically, our prior allows for the creation of abstract-time synthetic contexts, defined as collections of time series, similar enough to enable LaT-PFN to approximate the PPD.
+Specifically, our prior allows for the creation of synthetic contexts, defined as collections of time series, similar enough to enable LaT-PFN to approximate the PPD.
   • We demonstrate, through extensive experimentation across a variety of datasets, that LaT-PFN outperforms established baselines with limited training resources, on multiple tasks.
@@
 The PPD is typically defined as the probability distribution $P(y_{*}|x_{*},\mathcal{D})$, with $\mathcal{D}=\{(x_{i},y_{i})\}^{N}$ being a context dataset.
-The PFN framework works by approximating this distribution:
+PFNs work by approximating this distribution with network $Q_{\theta}(.)$:
 $Q_{\theta}(y_{*}|x_{*},\mathcal{D}_{j})\>$ $\approx\>P(y_{*}|x_{*},\mathcal{D})\quad\text{with}\quad\mathcal{D}_{j}\sim S(\mathcal{D}_{j}|\psi_{j}),\quad\psi_{j}\sim Pr(\psi)$
@@
 The network $Q$ is trained to perform meta-learning on any dataset $\mathcal{D}_{j}$ [41].
-The authors prove that this is a valid approximation to the PPD – see Appendix B.
+The authors prove that this is a valid approximation of the PPD – see Appendix C.
 Since its introduction, others have applied the PFN framework to various tasks [29, 16, 3, 22].
@@
 2.2 JEPA: Joint Embedding Predictive Architecture
-The Joint Embedding Predictive Architecture (JEPA) [35] framework aims to create strong embeddings of any modality, to efficiently handle complex relationships within data.
-The key concept is to predict state transitions $S_{t}\rightarrow S_{t+1}$ fully in latent space, in order to create embeddings that are not burdened by unnecessary detail.
+The Joint Embedding Predictive Architecture (JEPA) [35] framework aims to create a strong representation space to model complex relationships within data, without relying on reconstruction losses or heuristics.
+The key concept is to predict self-supervised state transitions $S_{t}\rightarrow S_{t+1}$ fully in latent space, yielding embeddings that are not burdened by unnecessary detail.
 Figure 2 illustrates the core components, which define a sequence of transformations aimed at processing inputs within latent space.
-Initially, an input embedder $\bar{x}=E_{x}(x)$ converts the input into latent space.
+Initially, an input embedding $\bar{x}=E_{x}(x)$ is produced by the input embedder.
 Figure 2: JEPA [35]
-This is followed by a predictor $\hat{\bar{y}}=P(\bar{x},z)$ which estimates the next latent state, using the transformed input and a variable $z$, intended to capture relationships not observable in the data space.
-A target embedder $\bar{y}=E_{y}(x,y)$ then maps the target, taking into account both $x$, $y$ and their the relationship.
+This is followed by a prediction $\hat{\bar{y}}=P(\bar{x},z)$ constituting the next estimated latent state, using the transformed input $\bar{x}$ and a variable $z$, intended to capture relationships not observable in data space – for example camera angles for computer vision.
+A target embedding $\bar{y}=E_{y}(x,y)$ is formed, taking into account both $x$, $y$ and their relationship.
 A criterion, typically a Mean Squared Error (MSE) loss $C(\hat{\bar{y}},\bar{y})$, evaluates the latent prediction.
-Finally, a generative decoder, $\hat{y}=D(\hat{\bar{y}})$, decodes the latent representations back to the data space.
+Finally, a generative decoder trained in isolation decodes the latent representations back to the data space $\hat{y}=D(\hat{\bar{y}})$.
 The JEPA framework has since been applied to image and video data [4, 5].
-The target encoder $y_{L}=E_{y}(x,y)$ is updated as an exponential moving average (EMA) of the input encoder $E_{x}$ and followed by a stop gradient, preventing collapse by learning an identity mapping [35, 5].
+The target encoder $y_{L}=E_{y}(x,y)$ is updated as an exponential moving average (EMA) of the input encoder $E_{x}$ and followed by a stop-gradient, preventing collapse by learning an identity mapping [35, 4].
 Other JEPA works include Motion [6], Audio [21], Point-cloud [46] and EEG transfer [27].
@@
 3.1 Problem Statement
-Figure 3: Problem statement of a time series forecasting PFN
+Figure 3: Our time series forecasting PFN problem statement, compared to preceding works [16, 29]
 Starting with a common definition of time series forecasting (TSF), we seek to derive a formulation of TSF for Bayesian Inference.
@@
 Where $\mathcal{D}$ is a context dataset with $N$ examples, for $\mathcal{D}\in\mathbb{R}^{N\times S\times F}$, sequence dimension $S$, example dimension $N$ and features $F$.
-This differs from the ForecastPFN formulation, which defines the context $\mathcal{D}\in\mathbb{R}^{S\times F}$ as simply the history of the target time series [17].
-This interpretation, although elegant, does not allow a context for the TSF task definition and is simply a regular forecasting task with simulated data.
+This differs from the ForecastPFN formulation, which defines the context $\mathcal{D}\in\mathbb{R}^{S\times F}$ as the history of the target time series [16].
+This interpretation, although elegant, does not allow for a context for the TSF task definition and is essentially a regular TSF approach trained with simulated data.
 See Figure 3 for a visualization.
-To complement this notation, let $\bar{x}=\bar{v}_{0:t}$ and latent target $\bar{y}=\bar{v}_{t+1:H}$ be the embedded versions of history $x=v_{0:t}$ and target $y=v_{t+1:H}$ respectively.
+To complement this notation, let embedded history $\bar{x}=\bar{v}_{0:t}$ and latent target $\bar{y}=\bar{v}_{t+1:H}$ be the embedded versions of history $x=v_{0:t}$ and target $y=v_{t+1:H}$ respectively.
 Similarly, let $\hat{\bar{y}}=\hat{\bar{v}}_{t+1:H}$ be the latent forecast, and $\hat{y}=\hat{v}_{t+1:H}$ be the forecast in data space.
-3.2 Abstract Time Dimension
-Unlike [17], instead of using absolute time units (year, month, etc.), we map time to a fixed abstract interval $T_{0}:T_{H}$ relative to the forecast start $t$ of each series, with a variable sampling rate – while ensuring the absence of any data leakage.
+3.2 Normalized Abstract Time-axis
+Unlike [16], instead of using absolute time units (year, month, etc.), we map time to a fixed normalized interval $T_{0}:T_{H}$ relative to the forecast start $t$ of each series, with a variable sampling rate – while ensuring the absence of any data leakage.
 This allows us to map context time series and past segments of the held-out series onto the same interval, independent of their actual place in history.
 The approach aligns well with the idea of using simulated data to learn Bayesian inference, as it creates a universal and consistent time dimension in which the model can more readily learn general temporal patterns.
-Furthermore, reducing the number of temporal degrees of freedom that require training, leads to a more efficient convergence and significantly decreases the computational demands for training.
+Furthermore, reducing the number of temporal degrees of freedom that require training, leads to a more efficient convergence and decreases the computational demands for training.
 3.3 An Architecture for Latent In-context Forecasting
@@
 These are fed into the PFN Predictor transformer, with the embedded held-out history prompts, using cross-attention.
-The latent predictions are compared to the latent target, then decoded with a stopgradient, and compared to real targets.
-Finally, we apply a supervised regularisation on simulation parameters.
+The latent predictions are compared to the latent target, then decoded with a stop-gradient, and compared to real targets.
+Finally, we apply a supervised regularization on simulation parameters.
 Embedder
 Figure 5: Production of embeddings and prompts by masking and selection
-The embedder $E(.)$ is an eight-layer dilated Mobilenet1D [30].
+The embedder $E_{\theta}$ is an eight-layer dilated Mobilenet1D [30].
 We empirically found convolutions to be superior to attention over the sequence dimension, in line with [63, 38].
@@
 First, we create the context embeddings $\bar{\mathcal{D}}$ and the latent target $\bar{y}$ with time $T$ and value $V$ features.
-For the latent target specifically, we use an EMA of the embedder parameters, apply a stop gradient, and select only the future timesteps:
+For the latent target specifically, we use an EMA of the embedder parameters, apply a stop-gradient, and select only the future timesteps:
 $\bar{\mathcal{D}}\hskip 5.0pt$ $=\hskip 5.0ptE_{\theta}(\mathcal{D})$
@@
 Predictor
-The predictor $PFN_{\theta}(.)$ forecasts the next latent state, given the prompt and context.
-We apply learned average pooling $AVG_{\theta}$ over the sequence dimension of the embedded context $\bar{\mathcal{D}}$, by cross-attention with a learned vector $\vec{q}\in\theta$.
-These are then provided to the encoder of the latent PFN predictor 4, with self-attention over the examples dimension.
-$\bar{\eta}\hskip 5.0pt=\hskip 5.0ptAVG_{\theta}(\bar{\mathcal{D}})$
+The predictor $\textit{PFN}_{\theta}$ forecasts the next latent state, given the prompt and context.
+We apply learned average pooling $\textit{AVG}_{\theta}$ over the sequence dimension of the embedded context $\bar{\mathcal{D}}$, by cross-attention with a learned vector $\vec{q}\in\theta$.
+These are then provided to the encoder of the latent PFN predictor in Figure 4, with self-attention over the examples dimension.
+$\bar{\eta}\hskip 5.0pt=\hskip 5.0pt\textit{AVG}_{\theta}(\bar{\mathcal{D}})$
 In order to apply 2D attention over the held-out series, we flatten the examples dimension $N$ and sequence dimension $S$ of the prompt $\bar{z}$.
-Next, we pass it to the decoder, whilst ensuring adjacent sequences from held-out series are not dependent on sharing information.
-$M\hskip 3.30002pt=\hskip 3.30002ptdiag(1)^{N\cdot S\times N\cdot S},\quad\hat{\bar{y}}\hskip 3.30002pt=\hskip 3.30002ptPFN_{\theta}(\bar{z},\bar{\eta},M),\quad\mathcal{L}_{\text{latent}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{\bar{y}}-\bar{y})^{2}$
+Next, we pass it to the decoder, whilst ensuring adjacent sequences from held-out series are not dependent on sharing information amongst each other.
+$M\hskip 3.30002pt=\hskip 3.30002ptdiag(1)^{N\cdot S\times N\cdot S},\quad\hat{\bar{y}}\hskip 3.30002pt=\hskip 3.30002pt\textit{PFN}_{\theta}(\bar{z},\bar{\eta},M),\quad\mathcal{L}_{\text{latent}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{\bar{y}}-\bar{y})^{2}$
 This is achieved by the diagonal-only mask $M$, which we adopt from the original PFN formulation [41]. Consequently, we stay true to the design from Figure 1.
 Decoder
-The decoder $D_{\theta}(.)$ is a three-layer feedforward network, trained with independent gradients in alignment with the JEPA methodology [35, 5, 6].
-We empirically found this to be critical for generating stable embeddings.
-The decoder utilizes a cross-entropy loss (see Appendix B) to define a categorical distribution over 100 segmented bins of the output space, incorporating label smoothing of 0.01 [50].
-This is made feasible by casting the values $V$ to an abstract space through z-normalizing with 2-std.
-By working in a value-abstract space, the decoder yields an output without making any assumptions on the target distribution – adhering closely to the PFN formulation [41].
-$\hat{y}\hskip 3.30002pt=\hskip 3.30002ptD_{\theta}(StopGrad(\hat{\bar{y}})),\quad\mathcal{L}_{\text{decoder}}\hskip 3.30002pt=\hskip 3.30002pt\mathbb{E}_{p(y)}\log p(\hat{y}\>|\>\>\hat{\bar{y}})$
+The decoder $D_{\theta}$ is a three-layer feedforward network tasked with decoding latent predictions, approximating the PPD in data space (see Appendix C).
+It is trained with independent gradients in alignment with the JEPA methodology [35, 4, 5].
+We empirically found this to be critical for generating stable embeddings upstream.
+The decoder utilizes a cross-entropy loss to define a categorical distribution over 100 segmented bins of the output space (see Appendix C), incorporating label smoothing of 0.01 [50].
+This is made feasible by z-normalizing the values $V$ with 2-std, towards a fixed interval.
+Consequently, the decoder yields an output without making any assumptions on the target distribution – adhering closely to the PFN formulation [41].
+[…]
 System Identification Head
@@
 This approach further supervises the latent space, directing it to the underlying characteristics of the simulation, such as trend and seasonality.
-This, in turn, leads to improved forecasting capabilities and stabilizes training [35].
-The latent embeddings and predictions are concatenated and averaged, then passed to a head $H_{\theta}(.)$ alongside the embedded context.
-$H_{\theta}(.)$ learns a multi-target regression towards unit-normalized simulation parameters.
-$\hat{y}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002ptH_{\theta}(AVG_{\theta}(\hat{\bar{y}},\bar{x}),\bar{C}),\quad\mathcal{L}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{y}_{\text{si}}-\psi_{j})^{2}$
+This, in turn, leads to improved forecasting capabilities and should stabilize training [35].
+The latent embeddings and predictions are concatenated and averaged, then passed to a head $H_{\theta}$ alongside the embedded context.
+$H_{\theta}$ learns a multi-target regression towards unit-normalized simulation parameters.
+$\hat{y}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002ptH_{\theta}\big{(}\>\textit{AVG}_{\theta}(\hat{\bar{y}},\bar{x}),\bar{C}\>\big{)},\quad\mathcal{L}_{\text{si}}\hskip 3.30002pt=\hskip 3.30002pt\frac{1}{n}\sum^{n}(\hat{y}_{\text{si}}-\psi_{j})^{2}$
 Loss
@@
 (f) Synthetic Prior
-Figure 6: Decoded fits plotted on the target, with history and target in blue and red scatter respectively, as well as the PPD fore- and backcast mean as blue line with variance, and finally the input example context in green in the background.
+Figure 6: Decoded fits plotted on the target, with history and target in blue and red scatter respectively, as well as the PPD fore- and backcast mean as a blue line with variance, and finally the input example context in green in the background.
 The fits demonstrate the predictions in latent space have picked up concepts such as seasonality and trend, making it easier to extrapolate after the forecast moment.
@@
 For this work, we adopted and adjusted ForecastPFN’s time series synthetic prior [16].
-Like ForecastPFN, we set hyperprior parameters that govern sampling simulation parameters.
+Like ForecastPFN, we set hyperprior parameters that govern the sampling of simulation parameters.
 Unlike ForecastPFN however, our context is defined not by the history of a single time series, but by a collection of example time series.
@@
   2. 2.
-    Context Parameters $(\alpha_{c})$: Ranges sampled uniformly from the hyperprior to define the limits within an individual context.
+    Context Parameters $(\alpha_{c})$: Ranges sampled uniformly from the hyperprior to define the limits of an individual context.
   3. 3.
-    Sub-context Parameters $(\mu)$: Cluster centers as $(\mu_{1},\mu_{2})$ sampled uniformly from the context range, are used to introduce contrast within a context and guide the model to recognize when a context example may be non-informative.
+    Sub-context Parameters $(\mu)$: Cluster centers $(\mu_{1},\mu_{2})$, sampled uniformly from the context range, are used to introduce contrast within a context and guide the model to recognize when a context example may be non-informative.
   4. 4.
-    Series Parameters $(\psi)$: Point estimates normally sampled using a fixed hyperprior variance $\sigma^{*}$, used as simulation parameters in $S(v_{i,0:T}|\psi_{i})$ to generate individual training data series.
+    Series Parameters $(\psi)$: Point estimates normally sampled with fixed hyperprior variance $\sigma^{*}$, used as simulation parameters in $S(v_{i,0:T}|\psi_{i})$ to generate individual training series.
 This approach ensures balance in inter- and intra-context variance, promoting coherence without trivializing the series and enhancing the model’s ability to generalize in a zero-shot setting.
-A full description of the synthetic prior and which parameters are used can be found in appendix C.1.
+A full description of the synthetic prior and which parameters are used can be found in Appendix D.1.
 4 Experimental Setup and Training Details
@@
 Additionally, we employ tuning on schedule parameters and $\mu$-parametrization, which was essential for training stability [59].
-For more training and tuning details, see Appendix C.
+For more training and tuning details, see Appendix D.
 Metrics
-We report the mean and error statistics of the zero-shot performance of multiple seeds.
+We report the mean and error statistics, assumed normally distributed, of the zero-shot performance of five seeds.
 These are accuracy, MSE per time-step, and the cumulative relative root mean squared error (CRRMSE).
 The latter is normalized by the cumulative target.
-We motivate choosing CRRMSE for being scale-invariant, and practical applicability in tasks such as sales forecasting.
+We motivate CRRMSE for its scale-invariance and practical applicability in tasks such as sales forecasting.
 Benchmarks & Context Curation
-For real datasets, fixed sequence length contexts were provided to the model.
-These were mapped to an abstract interval and Z-normalized with 2-std using the history.
-Which samples to include as contexts is coined Context Curation, and is a process analogous to Prompt Engineering in Large Language Models (LLMs) [40, 56].
-For details, see Appendix C.
+For real datasets, contexts were divided in fixed windows, normalized to a fixed time interval, and Z-normalized with 2-std using the history.
+The context-sample selection process is coined Context Curation, and is analogous to Prompt Engineering in Large Language Models (LLMs) [40, 56].
+For dataset-specific details see Appendix D.
 Baselines
@@
 Finally, all baselines are tuned, where applicable, and then evaluated on identical datasets.
-For more details on baselines, see Appendix C.
-Table 1: Forecasting scores (with std) for different time horizons. Results demonstrate a strong performance for all of the benchmarks from decoded latent predictions. For details see Appendix C
+For more details on baselines, see Appendix D.
+Table 1: Forecasting scores (with std) for different time horizons. Results demonstrate a strong performance for all of the benchmarks from decoded latent predictions. See data details in Appendix D
 Mean Squared Error (MSE)
@@
 FBProphet 0.61 0.98 0.67 0.96 0.29
-LaT-PFN 0.32 $\pm$ 4.0e-2 0.59 $\pm$ 3.0e-2 0.45 $\pm$ 1.8e-2 0.23 $\pm$ 2.3e-2 0.23 $\pm$ 2.8e-2
+LaT-PFN 0.32 $\pm$ 4.1e-2 0.59 $\pm$ 3.0e-2 0.45 $\pm$ 1.8e-2 0.23 $\pm$ 2.3e-2 0.23 $\pm$ 2.8e-2
 Cumulative Relative Root Mean Squared Error (CRRMSE)
@@
 FBProphet 5.75 8.33 6.42 7.79 3.34
-LaT-PFN 4.03 $\pm$ 2.4e-1 4.75 $\pm$ 2.1e-1 3.95 $\pm$ 1.5e-1 3.11 $\pm$ 1.8e-1 2.01 $\pm$ 2.0e-1
+LaT-PFN 4.03 $\pm$ 2.6e-1 4.75 $\pm$ 2.1e-1 3.67 $\pm$ 7.2e-1 3.11 $\pm$ 1.8e-1 2.01 $\pm$ 1.9e-1
 5 Results & Discussion
 Forecasting
-Table 2: Overall accuracy of UCR datasets, by fitting SVM on top of fixed-length summary embeddings [20]. Per-dataset scores are in Appendix E.
-training budget accuracy
-TS2Vec finetuned 45.7%
-LaT-PFN zero-shot 51.1%
+Table 2: Overall accuracy (with std) of all UCR datasets, by fitting SVM on top of fixed-length summary embeddings [19]. Per-dataset scores are in Appendix F.
+train budget accuracy
+TS2Vec 35 epochs 44.8% $\pm$ 1.1%
+LaT-PFN zero-shot 47.9% $\pm$ 3.7%
 We find that LaT-PFN exhibits strong, zero-shot forecasting performance across a wide variety of datasets, containing very different temporal patterns and levels of noise.
@@
 Next, we investigate the importance of the context by evaluating the uplift in performance on a synthetic validation set, for different context sizes.
-We chose to carry out this experiment on simulated data to guarantee a control environment.
+We chose to carry out this experiment on simulated data to guarantee a controlled environment.
 The result is visualized in Figure 8, which shows a clear correlation between performance and context size.
@@
 Classification
-We evaluate the quality of LaT-PFN’s embedding space by introducing a second downstream task, namely time series classification on the 129 multi-domain UCR datasets.
+We evaluate the quality of LaT-PFN’s embedding space by introducing a second downstream task, namely time series classification on the 128 multi-domain UCR datasets [19].
 We compare our model to TS2Vec, a state-of-the-art universal representation learning framework for time series [63].
@@
 We compare this to an SVM fitted on the fixed-length summary embeddings of a frozen LaT-PFN, trained on synthetic data only.
-As is evident in the results reported in Table 2, LaT-PFN outperforms TS2Vec across the entire benchmark of datasets.
-See Appendix E for individual dataset scores.
+As is evident in the results reported in Table 2, LaT-PFN outperforms TS2Vec on the combined datasets, despite the former being run in the zero-shot setting.
+See Appendix F for individual dataset scores.
 We argue these results may be due to bias introduced by time series-specific heuristics and contrastive positive pair selection, both of which TS2Vec and similar methods rely heavily upon [63, 60].
-By contrast, LaT-PFN independently discovers the most representative embedding space in which self-supervised learning is most optimal.
+By contrast, LaT-PFN independently discovers a representative embedding space in which self-supervised learning is highly efficient.
 Furthermore, TS2Vec is trained on the datasets it is evaluated on.
@@
 When analyzing the fixed-length summary embeddings after dimensionality reduction, we observe a clear separation in clusters by dataset – see Figure 7.
-This suggests that LaT-PFN picks up distributional differences between these domains, making these fixed-length summary embeddings useful for search, classification and other downstream applications.
-Figure 9: Embeddings per-timestep show patch-like behavior. More in Appendix D
-Interestingly we can observe the emergence of distinct, regular, and discrete patch-like tokens in the heatmaps of the embedding space.
+This suggests that LaT-PFN picks up distributional differences between these domains, making the fixed-length summary embeddings useful for search, classification, and other downstream applications.
+Figure 9: Normalized embeddings across sequence show patch-like behavior. More in Appendix E
+Interestingly, in the per-timestep embeddings we can observe the emergence of distinct, regular, and discrete patch-like tokens in the heatmaps of the embedding space.
 An example of this is visualized in Figure 9.
@@
 Finally, after applying dimensionality reduction to the set of patch embeddings (Figure 11), we notice the clear emergence of a few clusters.
-Given the aforementioned similarity between shapes of patches in data space with similar token embeddings, we speculate these could be a latent vocabulary that has been independently learned by the model, consisting of common "words" amongst multi-domain series.
+Considering the noted similarity between the shapes of patches in data space, and their corresponding averaged patch "token" embeddings, we hypothesize that these patches may represent a latent vocabulary independently learned by the model, comprising common "words" across multi-domain series.
 This would once more validate our core assumption that, despite time series data often being perceived as highly domain-dependant, there is in fact enough commonality to warrant the use of meta-learning approaches to create foundational models for this data modality.
@@
 Figure 11: DBScan [18] clustering of patches.
-For time series forecasting, focus has recently shifted from statistical models such as ARIMA [10], ES [25] and FBProphet [51] towards deep learning.
+For time series forecasting, the focus has recently shifted from statistical models such as ARIMA [9], ES [25] and FBProphet [51] towards deep learning.
 Early pioneers are DeepAR [47] and hybrid methods [49].
 Subsequently, N-Beats [42], N-hits [11], Informer [65] and FEDformer [66], combine classical with deep learning paradigms.
-More recently, following the success of the NLP domain, many attempts have been made to leverage the transformer architecture [54] to create foundation models [9] for time series forecasting.
+Following the success of the NLP domain, many attempts have been made to leverage the transformer architecture [54] to create foundation models [8] for time series forecasting.
 The two main approaches for doing so can be categorized as; (a) re-purposing LLMs [26, 67, 31] and (b) training transformer models from scratch [16, 3, 57, 14, 61, 45].
-By contrast, our research seeks to integrate representation- and in-context learning in time series forecasting, combining strong performance with a low computational budget.
+By contrast, our research seeks to integrate representation- and in-context learning in time series forecasting, combining strong performance with a low computational and data budget.
 Another approach altogether is to create embeddings and leave the downstream application to the discretion of end users.
 Notable are TS2Vec [63], TimeCLR [60], SimMTM [15] and TF-C [64].
-These methods often rely heavily on expert heuristics, whereas our approach serves both as an embedder and a forecaster, instead leveraging a self-supervised latent space to drive predictions.
+These methods often rely heavily on expert heuristics, whereas our approach serves both as an embedder and a forecaster, leveraging a self-supervised latent space to drive predictions.
 7 Limitations
@@
 This is beyond the scope of this paper but may be addressed in future work.
-Finally, we designed our model to be highly adaptive to a given context.
+Thirdly, we designed our model to be highly adaptive to a given context.
 As a result, sub-optimal contexts can lead to variability in performance.
@@
 In future work, we hope to address this by exploring automated methods – via retrieval augmented generation (RAG) [37] or prompt tuning [36].
+Finally, the model has proven somewhat sensitive to initialization and the normalization functions used.
 Addressing these points will be valuable for enhancing the model’s versatility.
@@
 LaT-PFN is trained exclusively on synthetic data, for which we define a novel prior.
-Combined with an abstract time dimension, this makes our model extremely versatile and able to achieve strong forecasting results across different data distributions, time granularities, and forecast horizons.
+Combined with a normalized abstract time-axis, this makes our model extremely versatile and able to achieve strong forecasting results across different data distributions, time granularities, and forecast horizons.
 Finally, its strong embedding space exhibits interesting emergent qualities, such as the development of patch-like tokens that resemble a corpus of time series features.
@@
 Appendix / supplemental material
-Appendix A Broader Societal Impacts
+Appendix A Code
+We provide the code for this work at https://github.com/StijnVerdenius/Lat-PFN.
+Appendix B Broader Societal Impacts
 It is our belief that the work presented in this paper does not pose major negative societal risks.
@@
 We hope this will result in widespread adoption of pre-trained models for forecasting tasks, thus improving industry effectiveness and reducing energy consumption and CO2 emissions.
-Appendix B Derivation PFN NLLLoss
-This derivation shows how to approximate Bayesian Inference via a PFN.
-It is adapted from [41] and forms the justification of why using a simulation with a cross-entropy loss is an approximation of the PPD.
-Assume we want to approximate PPD $P(y|x,\mathcal{D})$ with network $Q(y|x,\mathcal{D})$.
-We start with the KL-divergence between the two and end up at entropy function $\bm{H}()$ and constant $C$.
-The latter does not depend on parameters $\theta$ and is dropped in optimisation.
+Appendix C Derivation PFN NLLLoss
+This derivation, adapted from [41], illustrates the use of a cross-entropy loss with a synthetic prior as an approximation of the Predictive Posterior Distribution (PPD).
+This is a more effective way to learn Bayesian inference, as opposed to estimating the true posterior, which is often intractable.
+Furthermore, it yields a flexible approach by encoding expert data, which makes it highly adaptable for developing downstream applications [41].
+Assume we want to approximate PPD $P(y|x,\mathcal{D})$ with network $Q(y|x,\mathcal{D})$, for contexts $\mathcal{D}$.
+We start with the expectation over the contexts, of the KL-divergence between distributions $P(.)$ and $Q(.)$:
 […]
@@
 […]
-We can then integrate the PFN simulation by sampling the data from our prior:
+We find equivalence to the entropy function $\bm{H}(.)$ and constant $C$.
+The latter in independent of parameters $\theta$ and is dropped in optimization [41].
+Next, we elect to incorporate the PFN synthetic prior simulation by sampling data from our simulation:
 […]
 […]
-Appendix C Reproducibility Details
+These formulas show that, for the purpose of optimization, this is equivalent to applying the cross-entropy loss – a commonly used loss function.
+The output space is split in bins to allow the use of a cross-entropy loss for regression and forecasting problems – following the PFN example [41].
+Furthermore, the synthetic prior provides control over the prior distribution of contexts and access to infinite training data resources, including contexts, held-out examples, and targets [41].
+Finally, in this work, we introduce a separation of concerns between predicting and decoding.
+The end-to-end PPD task is effectively split into (a) learning a good way of summarizing the prior distribution within the latent prediction distribution $P(\hat{\bar{y}}|x,\mathcal{D},\theta)$, and then (b) separately learning an approximated posterior based on that summary $P(y|\hat{\bar{y}},\theta)$ – again without assumptions on output distribution family.
+Although the optimization is decoupled, there remains a causal relationship between the output distribution over $y$ and it dependent variables: $x$, $\mathcal{D}$, and parameters $\theta$.
+Therefore, we argue that this approach is a reasonable approximation of the PPD – given the latent predictions as an intermediary.
+Appendix D Reproducibility Details
 This section is dedicated to providing comprehensive reproducibility details for the research presented in this document.
 By including this appendix, we aim to uphold transparency standards in research.
-C.1 Exhaustive Description Synthesis Training Data
+D.1 Exhaustive Description Synthesis Training Data
 Following the example set by [16] we define a synthetic prior that leverages the underlying components of time series data, namely trend, seasonality, and noise.
@@
 $z\sim Weibull(1,k),~{}\overline{z}=\left(\ln{2}\right)^{1/k},~{}m_{noise}\in{\{\mathcal{U}(0,0.1),\mathcal{U}(0.2,0.4),\mathcal{U}(0.6,0.8)\}}$
-Unlike [17], we train using an abstract time dimension.
+Unlike [16], we train using an normalized time dimension.
 In order to include many different types of temporal patterns, we define a resolution parameter $\rho$, which determines how many points are represented per interval, for each individual time series.
@@
 $\delta_{\text{harmonics}}$ $4$ $12$
-C.2 Baselines
+D.2 Baselines
 Forecasting
@@
   • ARIMA [13], another classic baseline from the industry. We use the Statsmodel implementation [44] publicly available and pass the argument order=(5, 1, 0).
-  • ForecastPFN [17] is a flexible in-context forcaster model and one of the inspirations for this work.
+  • ForecastPFN [16] is a flexible in-context forecasting model and one of the inspirations for this work.
 We download their open-source repository available under Apache Licence 2.0 and trained model weights.
@@
 They open-source their training code under the MIT Licence, so we train TS2Vec from scratch on the train splits of the UCR [19] archive, after which we evaluate on the corresponding test splits with the resulting frozen TS2Vec architecture
-C.3 Benchmark Preprocessing
+D.3 Benchmark Preprocessing
 We use five datasets for our experiments, detailed in section 4. These are considered standard datasets for benchmarking the performance of time series forecasting models [43, 58, 65].
@@
 Additionally, we further split up the univariate series along the time dimension to create single entities of different granularity, which we then use for context creation.
-We normalize the time dimension of each dataset so that it fits in the abstract $[-3,1]$ interval.
+We normalize the time dimension of each dataset so that it fits within the normalized $[-3,1]$ interval.
 Furthermore, we apply Z-score normalization, with 2 standard deviations, to the values of each univariate time series.
@@
 However, practically we obtained the actual files from the ForecastPFN repository [16, 1]
-C.3.1 Illness
+D.3.1 Illness
 Illness is a dataset of influenza-like illness patients in the United States. It reports patients data in a weekly time granularity, from 1997 to 2024. [24]
@@
 We define a sequence length of 160 equidistant intervals, of which the last 25% are used as the prediction target. We use a rolling window over the dataset, with a stride of 1.
-We hand-picked 4 windows of the historic series as context examples and kept 1 as held-out starting from forecast date 01-01-2018, which we evaluate the zero-shot performance on.
+We hand-picked 8 windows of the historic series as context examples and kept 1 as held-out starting from forecast date 01-01-2018, which we evaluate the zero-shot performance on.
 Context Held-out
-ILITOTAL 2008-2012 ILITOTAL 2018-2022
-ILITOTAL 2013-2017
+AGE_65 2008-2012 ILITOTAL 2018-2022
+AGE_65 2013-2017
+AGE_0-4 2008-2012
+AGE_0-4 2013-2017
+AGE_5-24 2008-2012
+AGE_5-24 2013-2017
 TOTAL_PATIENTS 2008-2012
 TOTAL_PATIENTS 2013-2017
-C.3.2 EttH1 & EttH2
+The illness dataset is provided by a public government institution (CDC), which does not provide an explicit license.
+D.3.2 EttH1 & EttH2
 The ETT (Electricity Transformer Temperature) dataset is an electricity power deployment dataset.
@@
     ETTh1 MULL 2017/01 - 2017/05
-    ETTh1 OT 2017/01 - 2017/05
     ETTh2 HULL 2017/01 - 2017/05
@@
     ETTh2 MULL 2017/01 - 2017/05
-    ETTh2 OT 2017/01 - 2017/05
   • Experiment 2 - ETTh2:
@@
     ETTh2 HUFL 2017/01 - 2017/05
-    ETTh2 OT 2017/01 - 2017/05
-C.3.3 Traffic
+    ETTh2 HULL 2017/01 - 2017/05
+Both datasets are under Creative Commons Attribution-NoDerivatives 4.0 International.
+D.3.3 Traffic
 Traffic is a dataset reporting freeway congestion in California, USA.
@@
 405 2016/10 - 2017/04
-C.3.4 ECL
+D.3.4 ECL
 ECL (Electricity Consuming Load) [52] reports electricity consumption in Kwh with an hourly frequency from 2016-07-01 to 2019-07-02.
@@
 We define a sequence length of 240 equidistant intervals, of which the last 25% are used as targets. We use a rolling window over each time series with a stride of 10.
-We provide a context dimension of 15 and leave two held-out time series.
+We provide a context dimension of 30 and leave two held-out time series.
 Context Held-out
-OT 2019/01/01 - 2018/01/14 OT 2019/05/01 - 2018/05/14
-OT 2019/01/07 - 2018/01/20 OT 2019/04/01 - 2018/04/14
-OT 2019/01/17 - 2018/01/30
-OT 2019/02/01 - 2019/02/14
-OT 2019/02/07 - 2019/02/20
-OT 2019/02/17 - 2019/03/02
-OT 2019/03/01 - 2019/03/14
-OT 2019/03/07 - 2019/03/20
-OT 2019/03/17 - 2019/03/30
-OT 2017/04/01 - 2017/04/14
-OT 2017/04/07 - 2017/04/20
-OT 2017/04/17 - 2017/04/30
-C.4 Training & Tuning
+208 2018/04/01 - 2018/04/14 OT 2019/05/01 - 2018/05/14
+208 2018/04/07 - 2018/04/20 OT 2019/04/01 - 2018/04/14
+208 2018/04/17 - 2018/04/30
+208 2017/04/01 - 2017/04/14
+208 2017/04/07 - 2017/04/20
+208 2017/04/17 - 2017/04/30
+208 2019/01/01 - 2019/01/14
+208 2019/01/07 - 2019/01/20
+208 2019/01/17 - 2019/01/30
+208 2019/02/01 - 2019/02/14
+208 2019/02/07 - 2019/02/20
+208 2019/02/17 - 2019/03/02
+208 2019/03/01 - 2019/03/14
+208 2019/03/07 - 2019/03/20
+208 2019/03/17 - 2019/03/30
+313 2018/04/01 - 2018/04/14
+313 2018/04/07 - 2018/04/20
+313 2018/04/17 - 2018/04/30
+313 2017/04/01 - 2017/04/14
+313 2017/04/07 - 2017/04/20
+313 2017/04/17 - 2017/04/30
+313 2019/01/01 - 2019/01/14
+313 2019/01/07 - 2019/01/20
+313 2019/01/17 - 2019/01/30
+313 2019/02/01 - 2019/02/14
+313 2019/02/07 - 2019/02/20
+313 2019/02/17 - 2019/03/02
+313 2019/03/01 - 2019/03/14
+313 2019/03/07 - 2019/03/20
+313 2019/03/17 - 2019/03/30
+This dataset is under CC BY 4.0 licence.
+D.4 Training & Tuning
 Please consider the following training and tuning details:
@@
   • We train with a linear warmup of 95 epochs for JEPA-EMA decay and weight decay. Increasing the former from 0.9952 to 1.0 and the latter from 1.77e-4 to 4.9e-2.
-  • We tune all of the above hyperparameters with Optuna [3] in the 3 runs, spanning a total of 216 hours.
+  • We tune all of the above optimizer hyperparameters, $\lambda_{\text{di}}$ and $\lambda_{\text{latent}}$ with Optuna [2] in the 3 runs, spanning a total of 216 hours.
   • We apply MUP [59], which allows us to tune a smaller model and zero-shot transfer the hyperparameters we found.
@@
   • All the variables in the bullets below were not tuned but hand-picked by either trial and error or expert knowledge.
-  • We apply TF32 quantization but no mixed-precision training since this gave stability issues
+  • We apply TF32 quantization but no mixed-precision training since this gave stability issues.
   • For the final model we train with a batch size of 32, a context size of 14, a history of 180, horizon of 60 and held-out size of 2.
@@
   • Since we have a synthetic dataset, we have no natural end of an epoch. We elect to choose a 250-batch epoch before updating schedules.
-  • We map our training data to an abstract time domain of $[-3,1]$, with the forecast moment at the origin.
-  • We define 100 bins on the abstract $[-3.5,3.5]$ range, as the output space for the decoder.
+  • We map our training data to an normalized time domain of $[-3,1]$, with the forecast moment at the origin.
+  • We define 100 bins on the normalized $[-3.5,3.5]$ range, as the output space for the decoder.
   • Model-width is set at 512 neurons for the final model and 128 neurons for tuning. Corresponding depths are:
@@
 Figure 12: PCA plot of fixed-length summary embeddings grouped by domain
-Appendix D Latent Space Visualised
+Appendix E Latent Space Visualised
 (a) ETTh1
@@
 Figure 13: Additional per-step time series embeddings
-Appendix E Additional results
-E.1 UCR [20] Accuracy by Dataset
+Appendix F Additional results
+F.1 UCR [19] Accuracy by Dataset
 Table 4: UCR scores per dataset: A-F
 UCR Datasets Our work TS2Vec
-ACSF1_TEST 37.00% 41.00%
-Adiac_TEST 2.05% 2.05%
-AllGestureWiimoteX_TEST 29.14% 25.14%
-AllGestureWiimoteY_TEST 25.00% 17.86%
-AllGestureWiimoteZ_TEST 22.86% 16.43%
-ArrowHead_TEST 46.29% 41.71%
-BME_TEST 54.00% 46.67%
-Beef_TEST 50.00% 40.00%
-BeetleFly_TEST 60.00% 85.00%
-BirdChicken_TEST 60.00% 75.00%
-CBF_TEST 53.33% 47.56%
-Car_TEST 21.67% 21.67%
-Chinatown_TEST 60.06% 76.38%
-ChlorineConcentration_TEST 53.26% 53.26%
-CinCECGTorso_TEST 29.06% 25.58%
-Coffee_TEST 75.00% 82.14%
-Computers_TEST 56.40% 66.40%
-CricketX_TEST 29.23% 13.85%
-CricketY_TEST 24.87% 17.44%
-CricketZ_TEST 29.49% 14.36%
-Crop_TEST 38.66% 32.28%
-DiatomSizeReduction_TEST 30.07% 30.07%
-DistalPhalanxOutlineAgeGroup_TEST 67.63% 46.76%
-DistalPhalanxOutlineCorrect_TEST 58.33% 58.33%
-DistalPhalanxTW_TEST 58.27% 30.22%
-DodgerLoopDay_TEST 23.75% 37.50%
-DodgerLoopGame_TEST 56.52% 50.00%
-DodgerLoopWeekend_TEST 87.68% 89.13%
-ECG200_TEST 64.00% 64.00%
-ECG5000_TEST 82.69% 87.51%
-ECGFiveDays_TEST 49.71% 49.71%
-EOGHorizontalSignal_TEST 17.13% 29.83%
-EOGVerticalSignal_TEST 21.82% 20.99%
-Earthquakes_TEST 74.82% 74.82%
-ElectricDevices_TEST 51.87% 54.30%
-EthanolLevel_TEST 25.40% 26.40%
-FaceAll_TEST 26.63% 42.43%
-FaceFour_TEST 44.32% 18.18%
-FacesUCR_TEST 27.46% 18.29%
-FiftyWords_TEST 22.64% 12.53%
-Fish_TEST 14.29% 22.29%
-FordA_TEST 76.52% 61.52%
-FordB_TEST 65.43% 53.95%
-FreezerRegularTrain_TEST 76.70% 75.75%
-FreezerSmallTrain_TEST 76.00% 75.86%
-Fungi_TEST 60.22% 46.24%
+Acsf1 Test 36.00 % $\pm$ 5.0 % 43.00 % $\pm$ 2.8 %
+Adiac Test 03.07 % $\pm$ 2.0 % 02.05 % $\pm$ 0.0 %
+Allgesturewiimotex Test 24.50 % $\pm$ 5.8 % 21.21 % $\pm$ 5.6 %
+Allgesturewiimotey Test 24.39 % $\pm$ 5.0 % 16.36 % $\pm$ 2.1 %
+Allgesturewiimotez Test 19.25 % $\pm$ 4.9 % 17.50 % $\pm$ 1.5 %
+Arrowhead Test 43.00 % $\pm$ 10.2 % 44.86 % $\pm$ 4.4 %
+Bme Test 51.83 % $\pm$ 2.0 % 49.00 % $\pm$ 3.3 %
+Beef Test 41.67 % $\pm$ 8.8 % 38.33 % $\pm$ 2.4 %
+Beetlefly Test 56.25 % $\pm$ 4.8 % 75.00 % $\pm$ 14.1 %
+Birdchicken Test 67.50 % $\pm$ 15.5 % 70.00 % $\pm$ 7.1 %
+Cbf Test 47.92 % $\pm$ 10.8 % 40.33 % $\pm$ 10.2 %
+Car Test 21.67 % $\pm$ 0.0 % 21.67 % $\pm$ 0.0 %
+Chinatown Test 79.23 % $\pm$ 15.2 % 78.57 % $\pm$ 3.1 %
+Chlorineconcentration Test 53.65 % $\pm$ 0.6 % 53.26 % $\pm$ 0.0 %
+Cincecgtorso Test 25.85 % $\pm$ 2.1 % 25.18 % $\pm$ 0.6 %
+Coffee Test 72.32 % $\pm$ 5.4 % 82.14 % $\pm$ 0.0 %
+Computers Test 53.90 % $\pm$ 2.5 % 66.80 % $\pm$ 0.6 %
+Cricketx Test 21.15 % $\pm$ 8.7 % 13.85 % $\pm$ 0.0 %
+Crickety Test 21.47 % $\pm$ 10.9 % 16.54 % $\pm$ 1.3 %
+Cricketz Test 20.45 % $\pm$ 10.4 % 13.85 % $\pm$ 0.7 %
+Crop Test 31.78 % $\pm$ 6.2 % 32.07 % $\pm$ 0.3 %
+Diatomsizereduction Test 31.70 % $\pm$ 3.3 % 30.07 % $\pm$ 0.0 %
+Distalphalanxoutlineagegroup Test 63.49 % $\pm$ 8.3 % 46.76 % $\pm$ 0.0 %
+Distalphalanxoutlinecorrect Test 58.33 % $\pm$ 0.0 % 58.33 % $\pm$ 0.0 %
+Distalphalanxtw Test 59.35 % $\pm$ 3.7 % 30.22 % $\pm$ 0.0 %
+Dodgerloopday Test 26.88 % $\pm$ 5.2 % 38.12 % $\pm$ 0.9 %
+Dodgerloopgame Test 60.87 % $\pm$ 11.9 % 48.55 % $\pm$ 2.0 %
+Dodgerloopweekend Test 72.10 % $\pm$ 26.6 % 90.22 % $\pm$ 1.5 %
+Ecg200 Test 64.00 % $\pm$ 0.0 % 64.00 % $\pm$ 0.0 %
+Ecg5000 Test 86.03 % $\pm$ 4.0 % 86.33 % $\pm$ 1.7 %
+Ecgfivedays Test 52.41 % $\pm$ 5.4 % 49.71 % $\pm$ 0.0 %
+Eoghorizontalsignal Test 16.99 % $\pm$ 2.8 % 29.42 % $\pm$ 0.6 %
+Eogverticalsignal Test 16.44 % $\pm$ 5.2 % 21.13 % $\pm$ 0.2 %
+Earthquakes Test 74.82 % $\pm$ 0.0 % 74.82 % $\pm$ 0.0 %
+Electricdevices Test 52.58 % $\pm$ 3.0 % 52.41 % $\pm$ 2.7 %
+Ethanollevel Test 25.95 % $\pm$ 0.6 % 26.30 % $\pm$ 0.1 %
+Faceall Test 25.72 % $\pm$ 4.5 % 41.57 % $\pm$ 1.2 %
+Facefour Test 38.92 % $\pm$ 10.1 % 27.27 % $\pm$ 12.9 %
+Facesucr Test 23.33 % $\pm$ 3.8 % 16.32 % $\pm$ 2.8 %
+Fiftywords Test 22.64 % $\pm$ 5.6 % 12.53 % $\pm$ 0.0 %
+Fish Test 21.00 % $\pm$ 6.1 % 22.57 % $\pm$ 0.4 %
+Forda Test 67.42 % $\pm$ 10.8 % 62.88 % $\pm$ 1.9 %
+Fordb Test 57.69 % $\pm$ 6.4 % 56.11 % $\pm$ 3.1 %
+Freezerregulartrain Test 76.06 % $\pm$ 0.9 % 75.79 % $\pm$ 0.0 %
+Freezersmalltrain Test 73.27 % $\pm$ 5.2 % 75.86 % $\pm$ 0.0 %
+Fungi Test 67.74 % $\pm$ 10.2 % 46.24 % $\pm$ 0.0 %
 Table 5: UCR scores per dataset: G-R
 UCR Datasets Our work TS2Vec
-GestureMidAirD1_TEST 20.77% 20.00%
-GestureMidAirD2_TEST 26.92% 20.77%
-GestureMidAirD3_TEST 16.15% 10.00%
-GesturePebbleZ1_TEST 40.12% 50.58%
-GesturePebbleZ2_TEST 42.41% 31.01%
-GunPointAgeSpan_TEST 58.23% 57.91%
-GunPointMaleVersusFemale_TEST 62.66% 84.81%
-GunPointOldVersusYoung_TEST 52.38% 100.00%
-GunPoint_TEST 49.33% 64.67%
-Ham_TEST 59.05% 51.43%
-HandOutlines_TEST 64.05% 64.05%
-Haptics_TEST 22.40% 20.78%
-Herring_TEST 59.38% 59.38%
-HouseTwenty_TEST 69.75% 71.43%
-InlineSkate_TEST 18.18% 18.73%
-InsectEPGRegularTrain_TEST 100.00% 100.00%
-InsectEPGSmallTrain_TEST 100.00% 100.00%
-InsectWingbeatSound_TEST 30.15% 21.26%
-ItalyPowerDemand_TEST 61.22% 49.85%
-LargeKitchenAppliances_TEST 56.27% 57.60%
-Lightning2_TEST 54.10% 54.10%
-Lightning7_TEST 30.14% 41.10%
-Mallat_TEST 25.54% 12.32%
-Meat_TEST 68.33% 53.33%
-MedicalImages_TEST 51.45% 51.45%
-MelbournePedestrian_TEST 24.52% 62.03%
-MiddlePhalanxOutlineAgeGroup_TEST 45.45% 18.83%
-MiddlePhalanxOutlineCorrect_TEST 57.04% 57.04%
-MiddlePhalanxTW_TEST 46.10% 27.27%
-MixedShapesRegularTrain_TEST 59.13% 52.21%
-MixedShapesSmallTrain_TEST 48.78% 48.49%
-MoteStrain_TEST 87.22% 77.32%
-NonInvasiveFetalECGThorax1_TEST 21.63% 1.83%
-NonInvasiveFetalECGThorax2_TEST 26.97% 1.83%
-OSULeaf_TEST 35.95% 18.18%
-OliveOil_TEST 40.00% 40.00%
-Overall 51.20% 45.74%
-PLAID_TEST 31.10% 24.39%
-PhalangesOutlinesCorrect_TEST 61.31% 61.31%
-Phoneme_TEST 15.98% 11.29%
-PickupGestureWiimoteZ_TEST 38.00% 52.00%
-PigAirwayPressure_TEST 5.77% 19.71%
-PigArtPressure_TEST 4.81% 25.96%
-PigCVP_TEST 5.29% 46.63%
-Plane_TEST 22.86% 9.52%
-PowerCons_TEST 87.22% 85.56%
-ProximalPhalanxOutlineAgeGroup_TEST 82.44% 48.78%
-ProximalPhalanxOutlineCorrect_TEST 68.38% 68.38%
-ProximalPhalanxTW_TEST 63.90% 35.12%
-RefrigerationDevices_TEST 39.20% 54.40%
-Rock_TEST 46.00% 34.00%
+Gesturemidaird1 Test 25.77 % $\pm$ 6.6 % 23.08 % $\pm$ 4.4 %
+Gesturemidaird2 Test 30.00 % $\pm$ 7.2 % 21.15 % $\pm$ 0.5 %
+Gesturemidaird3 Test 14.81 % $\pm$ 6.9 % 08.46 % $\pm$ 2.2 %
+Gesturepebblez1 Test 25.44 % $\pm$ 11.9 % 45.64 % $\pm$ 7.0 %
+Gesturepebblez2 Test 27.37 % $\pm$ 13.6 % 26.27 % $\pm$ 6.7 %
+Gunpointagespan Test 52.53 % $\pm$ 3.8 % 54.43 % $\pm$ 4.9 %
+Gunpointmaleversusfemale Test 55.06 % $\pm$ 5.1 % 84.49 % $\pm$ 0.4 %
+Gunpointoldversusyoung Test 52.38 % $\pm$ 0.0 % 100.00 % $\pm$ 0.0 %
+Gunpoint Test 54.50 % $\pm$ 7.5 % 57.00 % $\pm$ 10.8 %
+Ham Test 58.81 % $\pm$ 6.7 % 51.43 % $\pm$ 0.0 %
+Handoutlines Test 64.05 % $\pm$ 0.0 % 64.05 % $\pm$ 0.0 %
+Haptics Test 21.19 % $\pm$ 0.8 % 20.78 % $\pm$ 0.0 %
+Herring Test 59.38 % $\pm$ 0.0 % 59.38 % $\pm$ 0.0 %
+Housetwenty Test 62.18 % $\pm$ 5.9 % 71.01 % $\pm$ 0.6 %
+Inlineskate Test 17.68 % $\pm$ 0.7 % 18.00 % $\pm$ 1.0 %
+Insectepgregulartrain Test 85.84 % $\pm$ 25.7 % 100.00 % $\pm$ 0.0 %
+Insectepgsmalltrain Test 76.00 % $\pm$ 24.5 % 100.00 % $\pm$ 0.0 %
+Insectwingbeatsound Test 34.20 % $\pm$ 8.1 % 18.89 % $\pm$ 3.4 %
+Italypowerdemand Test 62.05 % $\pm$ 17.6 % 49.85 % $\pm$ 0.0 %
+Largekitchenappliances Test 47.27 % $\pm$ 7.8 % 57.73 % $\pm$ 0.2 %
+Lightning2 Test 54.10 % $\pm$ 0.0 % 54.10 % $\pm$ 0.0 %
+Lightning7 Test 37.67 % $\pm$ 7.3 % 36.99 % $\pm$ 5.8 %
+Mallat Test 15.63 % $\pm$ 6.6 % 12.32 % $\pm$ 0.0 %
+Meat Test 67.50 % $\pm$ 1.7 % 55.83 % $\pm$ 3.5 %
+Medicalimages Test 51.51 % $\pm$ 0.1 % 51.45 % $\pm$ 0.0 %
+Melbournepedestrian Test 29.81 % $\pm$ 7.2 % 62.63 % $\pm$ 0.8 %
+Middlephalanxoutlineagegroup Test 34.09 % $\pm$ 17.9 % 18.83 % $\pm$ 0.0 %
+Middlephalanxoutlinecorrect Test 57.04 % $\pm$ 0.0 % 57.04 % $\pm$ 0.0 %
+Middlephalanxtw Test 38.15 % $\pm$ 12.0 % 27.27 % $\pm$ 0.0 %
+Mixedshapesregulartrain Test 48.98 % $\pm$ 14.4 % 50.85 % $\pm$ 1.9 %
+Mixedshapessmalltrain Test 44.85 % $\pm$ 7.1 % 49.20 % $\pm$ 1.0 %
+Motestrain Test 73.06 % $\pm$ 11.4 % 71.45 % $\pm$ 8.3 %
+Noninvasivefetalecgthorax1 Test 06.78 % $\pm$ 9.9 % 01.83 % $\pm$ 0.0 %
+Noninvasivefetalecgthorax2 Test 08.12 % $\pm$ 12.6 % 02.29 % $\pm$ 0.6 %
+Osuleaf Test 29.65 % $\pm$ 8.1 % 18.18 % $\pm$ 0.0 %
+Oliveoil Test 40.00 % $\pm$ 0.0 % 40.00 % $\pm$ 0.0 %
+Plaid Test 23.32 % $\pm$ 6.1 % 22.81 % $\pm$ 2.2 %
+Phalangesoutlinescorrect Test 61.31 % $\pm$ 0.0 % 61.31 % $\pm$ 0.0 %
+Phoneme Test 13.13 % $\pm$ 2.0 % 11.29 % $\pm$ 0.0 %
+Pickupgesturewiimotez Test 40.00 % $\pm$ 6.7 % 54.00 % $\pm$ 2.8 %
+Pigairwaypressure Test 06.61 % $\pm$ 1.7 % 19.71 % $\pm$ 0.0 %
+Pigartpressure Test 06.01 % $\pm$ 1.0 % 25.24 % $\pm$ 1.0 %
+Pigcvp Test 10.58 % $\pm$ 8.7 % 47.60 % $\pm$ 1.4 %
+Plane Test 26.67 % $\pm$ 16.7 % 09.52 % $\pm$ 0.0 %
+Powercons Test 88.06 % $\pm$ 3.2 % 85.00 % $\pm$ 0.8 %
+Proximalphalanxoutlineagegroup Test 66.22 % $\pm$ 19.9 % 48.78 % $\pm$ 0.0 %
+Proximalphalanxoutlinecorrect Test 68.38 % $\pm$ 0.0 % 68.38 % $\pm$ 0.0 %
+Proximalphalanxtw Test 50.49 % $\pm$ 17.8 % 35.12 % $\pm$ 0.0 %
+Refrigerationdevices Test 41.20 % $\pm$ 4.5 % 52.00 % $\pm$ 3.4 %
+Rock Test 44.50 % $\pm$ 3.4 % 39.00 % $\pm$ 7.1 %
 Table 6: UCR scores per dataset: S-Z
 UCR Datasets Our work TS2Vec
-ScreenType_TEST 45.33% 40.80%
-SemgHandGenderCh2_TEST 74.33% 58.83%
-SemgHandMovementCh2_TEST 30.89% 35.11%
-SemgHandSubjectCh2_TEST 37.11% 35.56%
-ShakeGestureWiimoteZ_TEST 54.00% 64.00%
-ShapeletSim_TEST 50.56% 87.78%
-ShapesAll_TEST 28.00% 42.67%
-SmallKitchenAppliances_TEST 40.53% 55.47%
-SmoothSubspace_TEST 57.33% 78.00%
-SonyAIBORobotSurface1_TEST 59.73% 42.93%
-SonyAIBORobotSurface2_TEST 78.07% 61.70%
-StarLightCurves_TEST 82.93% 84.74%
-Strawberry_TEST 64.32% 64.32%
-SwedishLeaf_TEST 31.68% 8.48%
-Symbols_TEST 55.78% 17.39%
-SyntheticControl_TEST 50.33% 77.33%
-ToeSegmentation1_TEST 70.61% 65.79%
-ToeSegmentation2_TEST 80.77% 80.77%
-Trace_TEST 65.00% 19.00%
-TwoLeadECG_TEST 49.96% 49.96%
-TwoPatterns_TEST 31.62% 25.87%
-UMD_TEST 49.31% 50.00%
-UWaveGestureLibraryAll_TEST 51.84% 12.37%
-UWaveGestureLibraryX_TEST 57.31% 22.28%
-UWaveGestureLibraryY_TEST 44.89% 20.38%
-UWaveGestureLibraryZ_TEST 52.07% 31.38%
-Wafer_TEST 89.21% 89.21%
-Wine_TEST 50.00% 50.00%
-WordSynonyms_TEST 29.31% 21.94%
-WormsTwoClass_TEST 62.34% 57.14%
-Worms_TEST 46.75% 42.86%
-Yoga_TEST 52.97% 53.57%
+Screentype Test 37.33 % $\pm$ 5.4 % 41.33 % $\pm$ 0.8 %
+Semghandgenderch2 Test 68.29 % $\pm$ 5.8 % 55.25 % $\pm$ 5.1 %
+Semghandmovementch2 Test 27.44 % $\pm$ 7.0 % 32.00 % $\pm$ 4.4 %
+Semghandsubjectch2 Test 40.67 % $\pm$ 8.7 % 33.67 % $\pm$ 2.7 %
+Shakegesturewiimotez Test 46.00 % $\pm$ 6.3 % 71.00 % $\pm$ 9.9 %
+Shapeletsim Test 60.83 % $\pm$ 8.7 % 83.33 % $\pm$ 6.3 %
+Shapesall Test 29.46 % $\pm$ 1.0 % 43.00 % $\pm$ 0.5 %
+Smallkitchenappliances Test 48.60 % $\pm$ 9.5 % 55.33 % $\pm$ 0.2 %
+Smoothsubspace Test 61.17 % $\pm$ 5.3 % 77.33 % $\pm$ 0.9 %
+Sonyaiborobotsurface1 Test 47.67 % $\pm$ 8.1 % 42.93 % $\pm$ 0.0 %
+Sonyaiborobotsurface2 Test 67.24 % $\pm$ 7.7 % 61.70 % $\pm$ 0.0 %
+Starlightcurves Test 76.83 % $\pm$ 12.8 % 83.96 % $\pm$ 1.1 %
+Strawberry Test 64.32 % $\pm$ 0.0 % 64.32 % $\pm$ 0.0 %
+Swedishleaf Test 25.72 % $\pm$ 6.0 % 07.84 % $\pm$ 0.9 %
+Symbols Test 43.52 % $\pm$ 20.0 % 17.39 % $\pm$ 0.0 %
+Syntheticcontrol Test 62.67 % $\pm$ 14.2 % 82.17 % $\pm$ 6.8 %
+Toesegmentation1 Test 62.17 % $\pm$ 5.9 % 66.89 % $\pm$ 1.6 %
+Toesegmentation2 Test 75.77 % $\pm$ 11.8 % 73.08 % $\pm$ 10.9 %
+Trace Test 52.25 % $\pm$ 11.1 % 19.00 % $\pm$ 0.0 %
+Twoleadecg Test 51.51 % $\pm$ 3.1 % 49.96 % $\pm$ 0.0 %
+Twopatterns Test 34.88 % $\pm$ 4.1 % 25.87 % $\pm$ 0.0 %
+Umd Test 52.78 % $\pm$ 5.9 % 55.90 % $\pm$ 8.3 %
+Uwavegesturelibraryall Test 26.95 % $\pm$ 18.5 % 12.33 % $\pm$ 0.1 %
+Uwavegesturelibraryx Test 57.49 % $\pm$ 4.6 % 17.23 % $\pm$ 7.1 %
+Uwavegesturelibraryy Test 48.75 % $\pm$ 6.1 % 19.11 % $\pm$ 1.8 %
+Uwavegesturelibraryz Test 51.70 % $\pm$ 5.4 % 25.49 % $\pm$ 8.3 %
+Wafer Test 90.97 % $\pm$ 3.5 % 89.21 % $\pm$ 0.0 %
+Wine Test 50.00 % $\pm$ 0.0 % 50.00 % $\pm$ 0.0 %
+Wordsynonyms Test 27.35 % $\pm$ 4.5 % 21.94 % $\pm$ 0.0 %
+Wormstwoclass Test 58.44 % $\pm$ 2.6 % 57.14 % $\pm$ 0.0 %
+Worms Test 43.83 % $\pm$ 1.9 % 42.86 % $\pm$ 0.0 %
+Yoga Test 53.34 % $\pm$ 0.3 % 53.57 % $\pm$ 0.0 %
 End of the Appendix.